### v1

# Sparse Estimation of Inverse Covariance and Partial Correlation Matrices via Joint Partial Regression

Abstract
We present a new method for estimating high-dimensional sparse partial correlation and inverse covariance matrices, which exploits the connection between the inverse covariance matrix and linear regression. The method is a two-stage estimation method wherein each individual feature is regressed on all other features while positive semi-definiteness is enforced simultaneously. We provide statistical rates of convergence for the proposed method which match, and improve upon, the state-of-the-art for inverse covariance and partial correlation matrix estimation, respectively. We also propose an efficient proximal splitting algorithm for numerically computing the estimate. The effectiveness of the proposed method is demonstrated on both synthetic and real-world data.
1 Introduction
Two important and closely related problems in statistical learning are the problems of estimating a partial correlation network and the inverse covariance matrix, also known as the precision matrix, from data. Partial correlation networks, which generalize the Gaussian graphical model, are used to model the relationships between variables while conditioning on all other variables, and are useful for inferring causal relationships between variables. Partial correlation networks are used in a plethora of applications, such as in the analysis of gene expression data, where the goal is to infer the regulatory relationships between genes (de la Fuente et al., 2004), and psychological data, where networks are used to model the relationships between psychological variables such as mood and attitude (Epskamp and Fried, 2018). The precision matrix, from which we can obtain the partial correlation network, is also of interest in its own right, as it also appears in linear discriminant analysis (Hastie et al., 2009) and in Markowitz portfolio selection (Markowitz, 1952). However, due to the high-dimensionality of the problem, estimating a precision or partial correlation matrix is often challenging as the number of parameters are on the order of the squared number of features. For this reason, classical methods such as using the inverse of the sample covariance matrix, are known to perform poorly whenever the number of observation is not extremely large. Additionally they produce estimates which are almost surely dense. This makes regularization crucial, since in many applications we typically only have a moderate number of observations, and in particular we are most often seeking a sparse estimate which gives rise to a more parsimonious and interpretable network model.
Related work.
Gaussian graphical models, and by extension sparse precision matrix estimation, have been studied extensively in the literature with the covariance selection problem originally being due to Dempster (1972). Today there are several methods for estimating the precision matrix, which can be divided into three main categories: methods based upon maximum likelihood estimation, methods based upon approximate inversion, and methods based upon what we shall refer to as partial regression. T/he partial regression approach exploits a connection between the precision matrix and linear regression, and is given its name due to its close relation to partial correlations.
Perhaps the most popular method of estimation is $\ell_{1}$-regularized maximum likelihood estimation (commonly known as the graphical lasso). This method was introduced independently by Yuan and Lin (2007) and d’Aspremont et al. (2008) (having slight variations), where the former work also provided a consistency result in the high-dimensional setting. There has additionally been much work on the problem of numerically solving the graphical lasso problem, such as the work of Friedman et al. (2008), who also coined the name.
Cai et al. (2011) provide a method of the second kind, based on minimizing the $\ell_{1}$-norm of the estimate under the constraint that the estimate is sufficiently close to an inverse of the sample covariance matrix in the $\ell_{\infty}$-norm. Robust extensions of this method are provided by Wang and Gu (2017) and Zhao and Liu (2014), with the former focusing on the case of contaminated data and the latter on the case of data drawn from a heavy-tailed distribution.
The present work is most closely related to the works of Meinshausen and Bühlmann (2006), vande Geer et al. (2014) and Yuan (2010), who opt for estimation via partial regression. Meinshausen and Bühlmann specifically consider the sparsity pattern estimation problem in a high-dimensional setting, and use the lasso (Tibshirani, 1996) to estimate the neighborhood of each node in the graphical model. Their work provides consistency results of their proposed method, and has become one of the most cited works on the subject. The work of vande Geer et al. continues in the same vein and estimate the precision matrix with the purpose of constructing confidence regions for high-dimensional models. Yuan provides a method that can be boiled down to solving a sequence of linear programs, opting for the Dantzig selector (Candès and Tao, 2007) in lieu of the lasso, and projecting the estimate onto the space of symmetric matrices in the operator $\ell_{1}$-norm. The author also proves that their method achieves the minimax optimal rate in the operator $\ell_{1}$-norm over certain parameter spaces.
Some work has additionally been done on the problem of estimating high-dimensional partial correlation matrices. Peng et al. (2009) provide a non-convex method dubbed SPACE for estimating the partial correlation matrix via partial regression. Friedman et al. (2010) and Khare et al. (2015) also propose pseudo-likelihood methods based on partial regression. Peng et al. and Khare et al. provide consistency results for their methods, with identical statistical rates of convergence.
Contributions.
The main contributions of the present work are three-fold, and are as follows:
  • We propose a novel method for simultaneously estimating a sparse precision and partial correlation matrix via linear regression, which we call joint partial regression. We also provide a robust extension of the method that uses sparse Huber regression, which is well-suited for contaminated data. For sub-Gaussian design, we provide statistical estimation error rates in a high-dimensional setting, which match the best known rates for precision matrix estimation and improve upon the best known rates for partial correlation matrix estimation.
  • We provide an efficient proximal algorithm for numerically fitting the proposed method which is well-suited for large-scale data sets. An implementation of the algorithm is also given in the form of a Python package written in Rust, which is available at https://github.com/samericks/joint-partial-regression.
  • We conduct numerical studies on synthetic data, showing that the proposed method outperforms the graphical lasso in terms of estimation error. We also apply the method to real stock market data, showing that the method is capable of capturing the underlying structure of the data.
Notation.
We write $\mbox{\bf S}^{p}$ for the set of real $p\times p$ matrices, and for $\Omega\in\mbox{\bf S}^{p}$ we write $\Omega\succeq 0$ if $\Omega$ is positive semi-definite. If $\Omega_{1}$ and $\Omega_{2}$ are both in $\mbox{\bf S}^{p}$ we write $\Omega_{1}\succeq\Omega_{2}$ if $\Omega_{1}-\Omega_{2}\succeq 0$. We write $\|\Omega\|_{\ell_{q}}=\max_{\|\nu\|_{q}\leq 1}\|\Omega\nu\|_{q}$ for the operator $\ell_{q}$-norm, and the Frobenius norm is defined as $\|\Omega\|_{\rm F}=\sqrt{\mathop{\bf Tr}(\Omega^{2})}$. The cardinality function $\mathop{\bf card}$ gives the number of elements if given a set argument, and the number of non-zero entries if given a vector argument.
2 The precision matrix and partial correlation
Likely the most common use of the precision matrix is in the context of Gaussian graphical models, where the connection between the precision matrix and the conditional independence structure of jointly Gaussian random variables is exploited. The basis of this connection is in fact a connection that the precision matrix also has to linear regression. Suppose $Z\in{\mbox{\bf R}}^{p}$ is a square-integrable mean-zero random vector with non-singular covariance $\Sigma$ and inverse $\Omega=\Sigma^{-1}$. Then the best linear unbiased predictor of a component $Z_{j}$ given all other components
$Z_{\setminus j}=(Z_{1},\dots,Z_{j-1},Z_{j+1},\dots,Z_{p})$
is $Z_{\setminus j}^{\top}\theta_{j}$, where
$\theta_{j}=\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j}.$
Here $\Sigma_{\setminus j,\setminus j}$ is the sub-matrix of $\Sigma$ obtained by removing the $j$-th row and column, and $\Sigma_{\setminus j,j}$ is the $j$-th column of $\Sigma$ with the $j$-th element removed. Likewise, the residual $\varepsilon_{j}=Z_{j}-Z_{\setminus j}^{\top}\theta_{j}$ has mean zero and variance $\tau_{j}^{2}=\Sigma_{jj}-\Sigma_{j,\setminus j}\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j}$. Using Schur complements, we have that
$\Omega_{jj}$ $=(\Sigma_{jj}-\Sigma_{j,\setminus j}\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j})^{-1},$
$\Omega_{\setminus j,j}$ $=-(\Sigma_{jj}-\Sigma_{j,\setminus j}\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j})^{-1}\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j},$
and thus
$\Omega_{jj}$ $=1/\tau_{j}^{2},$
$\Omega_{\setminus j,j}$ $=-\theta_{j}/\tau_{j}^{2}.$
In the Gaussian case we find that $\varepsilon_{j}$ is independent of $Z_{\setminus j}$, which then implies that $Z_{j}$ is conditionally independent of $Z_{k}$, $j\neq k$, given $Z_{\setminus\{j,k\}}$ if and only if $\theta_{jk}=0$, i.e., $\Omega_{jk}=0$. More generally, the partial correlation coefficient between $Z_{j}$ and $Z_{k}$ given $Z_{\setminus\{j,k\}}$ is given by the negative correlation between $\varepsilon_{j}$ and $\varepsilon_{k}$,
$\rho_{jk\mid\setminus\{j,k\}}=-\mathop{\bf Corr}(\varepsilon_{j},\varepsilon_{k})=-\frac{{\mathbb{E}}(\varepsilon_{j}\varepsilon_{k})}{\sqrt{{\mathbb{E}}(\varepsilon_{j}^{2}){\mathbb{E}}(\varepsilon_{k}^{2})}},$
which can be written in terms of the precision matrix and the linear regression as
$\rho_{jk\mid\setminus\{j,k\}}=-\frac{\Omega_{jk}}{\sqrt{\Omega_{jj}\Omega_{kk}}}=\frac{\tau_{k}}{\tau_{j}}\theta_{jk},$
respectively. Thus we define the partial correlation matrix $Q$ as
$Q=-T\Omega T,$
where $T=\mathop{\bf diag}(\tau_{1},\dots,\tau_{p})$. Note that some works (Schäfer and Strimmer, 2004; Marrelec et al., 2006) define the partial correlation matrix as $-T\Omega T+2I$, so that the diagonal elements are ones.
3 Proposed method
In light of the connection between the precision matrix and linear regression described in §2, a way to estimate the precision matrix is to form estimates of the linear regression coefficients $\hat{\theta}_{j}$ and the residual variances $\hat{\tau}_{j}^{2}$ for each $j$ and then compute an estimate of the precision matrix $\widehat{\Omega}$ via
$\widehat{\Omega}_{jj}$ $=1/\hat{\tau}_{j}^{2},$
$\widehat{\Omega}_{\setminus j,j}$ $=-\hat{\theta}_{j}/\hat{\tau}_{j}^{2},$
and an estimate of the partial correlation matrix $\widehat{Q}$ via
$\widehat{Q}=-\widehat{T}\widehat{\Omega}\widehat{T},$
where $\widehat{T}=\mathop{\bf diag}(\hat{\tau}_{1},\dots,\hat{\tau}_{p})$. An obvious approach, since we are interested in sparsity, is to use the lasso. Suppose we have access to a design matrix $X\in{\mbox{\bf R}}^{n\times p}$, where each row contains a sample of the random vector $Z$ and each column corresponds to a feature of the data. We can then estimate the linear regression coefficients by solving the lasso problem
$\hat{\theta}_{j}=\operatornamewithlimits{argmin}_{\theta\in\mathbf{R}^{p-1}}\left\{\frac{1}{2n}\|X_{j}-X_{\setminus j}\theta\|_{2}^{2}+\lambda\|\theta\|_{1}\right\}$
and subsequently estimate the residual variances by $\hat{\tau}_{j}^{2}=(1/n)\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}$. However, this approach does not result in a positive semi-definite estimate of the precision matrix, let alone a symmetric one. If we are solely interested in the unweighted partial correlation network, we could choose to include the edge $(j,k)$ only if $\widehat{\Omega}_{jk}$ and $\widehat{\Omega}_{kj}$ are both non-zero or, alternatively, if either is non-zero. If we are interested in the partial correlations themselves we could choose to symmetrize the estimate by averaging the estimates of $\widehat{\Omega}_{jk}$ and $\widehat{\Omega}_{kj}$, or by projection onto the cone of positive semi-definite matrices in some norm. We will however take a different approach, in which we will enforce positive semi-definiteness while simultaneously estimating the linear regression coefficients, given initial estimates of the residual variances.
3.1 Joint partial regression
We propose the following convex program that regresses the features at the same time as ensuring positive semi-definiteness,
$\begin{array}[]{ll}\text{minimize}&\displaystyle\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}-X_{\setminus j}\theta_{j}\|_{2}^{2}+\lambda\|\theta_{j}\|_{1}\right)\\ \text{subject to}&\Omega_{jj}=1/\hat{\tau}_{j}^{2},\quad\Omega_{\setminus j,j}=-\theta_{j}/\hat{\tau}_{j}^{2},\quad j=1,\dots,p,\\ &\Omega\succeq 0,\quad Q=-\widehat{T}\Omega\widehat{T}.\end{array}$ (1)
Here $\Omega\in{\mathbf{R}}^{p\times p}$ and $\theta_{j}\in{\mathbf{R}}^{p-1}$, $j=1,2,\dots,p$, are the optimization variables, and the data are the data matrix $X$, the estimates $\hat{\tau}_{j}^{2}$ of the residual variances, and the regularization parameter $\lambda\geq 0$. Note that the equality $Q=-\widehat{T}\Omega\widehat{T}$ is not a real constraint as $Q$ does not appear in the objective function or other constraints, but simply expresses that $Q$ is a by-product of the optimization problem in the stated way. We call the solutions $\widehat{\Omega}$ and $\widehat{Q}$ the joint partial regression estimates of the precision matrix and the partial correlation matrix, respectively. In order to obtain the initial estimates of the residual variances we can use the aforementioned lasso estimates of the linear regression coefficients to compute the sample variances $\hat{\tau}_{j}^{2}=(1/n)\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}$. We summarize the proposed procedure in Algorithm 3.1.
Note moreover that applying the derivation in §2 to the sample covariance matrix reveals that Algorithm 3.1 exactly retrieves the inverse sample covariance matrix when $\lambda=0$.
Input: Data matrix $X\in{\mbox{\bf R}}^{n\times p}$, penalty parameter $\lambda$.
for $j=1,\dots,p$ do
lasso regression
$\hat{\theta}_{j}=\operatornamewithlimits{argmin}_{\theta\in{\mathbf{R}}^{p-1}}\left\{\frac{1}{2n}\|X_{j}-X_{\setminus j}\theta\|_{2}^{2}+\lambda\|\theta\|_{1}\right\},$
and compute the estimate $\hat{\tau}_{j}^{2}=(1/n)\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}$ of the residual variance.
end
Solve the convex program (1) with the estimated residual variances $\hat{\tau}_{j}^{2}$ and regularization parameter $\lambda$ to obtain the estimates $\widehat{\Omega}$ and $\widehat{Q}$ of the precision matrix and partial correlation matrix, respectively.
Algorithm 3.1 Joint partial regression
3.2 Extensions
A benefit of our proposed method is that the convex program (1) can easily be modified. For example, we need not be limited to using quadratic loss, and may opt for another convex loss $\ell\colon{\mbox{\bf R}}^{n}\to{\mbox{\bf R}}$. In the case of contaminated or heavy-tailed data, we can benefit from taking $\ell$ as the Huber loss,
$\ell(z)=\frac{1}{n}\sum_{i=1}^{n}\phi_{\rho}(z_{i}),\quad\phi_{\rho}(r)=\begin{cases}r^{2}/2,&\text{if }|r|\leq\rho,\\ \rho(|r|-\rho/2),&\text{if }|r|>\rho.\end{cases}$ (2)
We may also use different regularization parameters $\lambda_{1},\dots,\lambda_{p}$ for each linear regression, rather than just a single $\lambda$. In practice, choosing these parameters via cross-validation does in fact not become computationally infeasible, since in the initial regression step we may simply perform $p$ model selection procedures for the separate lasso regressions in order to choose the parameters, and subsequently solve the main optimization problem only once.
Another possible modification to (1) is to constrain the eigenvalues of the precision matrix estimates to lie within $[\alpha,\beta]$ for some choice of $\alpha$ and $\beta$ with $0\leq\alpha\leq\beta\leq\infty$. In this way we can control the condition number of our estimate, which is desirable if we would like to produce an estimate of the covariance matrix via inversion. We thus have the generalized joint partial regression problem
$\begin{array}[]{ll}\text{minimize}&\displaystyle\sum_{j=1}^{p}\left(\ell(X_{j}-X_{\setminus j}\theta_{j})+\lambda_{j}\|\theta_{j}\|_{1}\right)\\ \text{subject to}&\Omega_{jj}=1/\hat{\tau}_{j}^{2},\quad\Omega_{\setminus j,j}=-\theta_{j}/\hat{\tau}_{j}^{2},\quad j=1,\dots,p\\ &\alpha I\preceq\Omega\preceq\beta I,\quad Q=-\widehat{T}\Omega\widehat{T}.\end{array}$ (3)
We can also choose another method than sample variance for estimating the residual variances. For example, Fan et al. (2012) suggested
$\hat{\tau}_{j}^{2}=\frac{1}{n-\hat{s}_{j}}\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2},\quad\hat{s}_{j}=\mathop{\bf card}(\hat{\theta}_{j}),$
which may be more suitable in the $p\gg n$ case. Under the contaminated data assumption we propose using Huber loss and a simple truncated mean squared estimator, where the residuals with magnitude larger than the threshold $\rho$ (the residuals that the Huber regression “rejected”) are left out. That is, for the residual $r_{j}=X_{j}-X_{\setminus j}\hat{\theta}_{j}$ we estimate the residual variance as
$\hat{\tau}_{j}^{2}=\frac{1}{n_{j}}\sum_{i=1}^{n}r_{ji}^{2}\mathbf{1}(|r_{ji}|\leq\rho),$ (4)
where $n_{j}=\sum_{j=1}^{p}\mathbf{1}(|r_{ji}|\leq\rho)$ is the number of residuals not omitted.
4 Theoretical results
In this section we provide theoretical guarantees for the estimation error of the proposed method that hold with high probability under given assumptions. Each data point, a random vector in ${\mbox{\bf R}}^{p}$, is assumed to have a true, unknown, covariance matrix that we denote by $\Sigma^{\star}$. The corresponding true precision matrix is $\Omega^{\star}=(\Sigma^{\star})^{-1}$. Our first assumption concerns how the dimensionality of the problem grows. Indeed, the theoretical results below are asymptotic in nature, assuming that both $n$ and $p$ tend to infinity. This first assumption specifies how fast $p$ can grow with $n$, and also bounds the sparsity of the true precision matrix.
Assumption 4.1
The dimensionality is such that $p/n\leq 1-\delta$ for some $\delta\in(0,1)$, and the degree
$d=\max_{j}\sum_{k\neq j}\mathbf{1}(\Omega_{jk}^{\star}\neq 0)$
of the partial correlation network is such that $d\sqrt{\log(p)/n}\leq M$ for some constant $M>0$.
Notice that the dependence on $n$ and $p$ is in many cases not made explicit; for example, in the above assumption both $\Omega^{\star}$ and $d$ tacitly depend on $n$ and $p$. We remark that the term $d\sqrt{\log(p)/n}$ that appears in the assumption has been shown by Yuan (2010) to be the minimax optimal error rate in the operator $\ell_{1}$-norm for precision matrix estimation. It is also the error rate of the graphical lasso in the spectral norm as shown by Ravikumar et al. (2011). Hence the boundedness assumption on that term is comparatively weak.
Our second assumption is that the data is sub-Gaussian. A real-valued random variable $Y$ is said to be sub-Gaussian if there exists a constant $C$ such that the tail probabilities satisfy
${\mathbb{P}}(|Y|\geq t)\leq 2\exp(-t^{2}/C^{2}),\quad\text{ for all }t\geq 0.$
For sub-Gaussian random variables we also define the sub-Gaussian norm $\|\cdot\|_{\psi_{2}}$ by
$\|Y\|_{\psi_{2}}=\inf\{t>0\colon{\mathbb{E}}(\exp(Y^{2}/t^{2}))\leq 2\}.$
With this in mind, we state the following assumption.
Assumption 4.2
The rows of the design matrix $X\in{\mbox{\bf R}}^{n\times p}$ are $n$ i.i.d. samples from a random vector with covariance matrix $\Sigma^{\star}$, and each $X_{ij}$ is sub-Gaussian with associated norm $\|X_{ij}\|_{\psi_{2}}\leq K$ for some $K>0$.
Since the bound $K$ is for all $p$, $i$ and $j$, we may say that the elements of $X$ are uniformly sub-Gaussian.
Finally we assume that the eigenvalues of true covariance matrix do not degenerate to zero or infinity, and that the maximum absolute row sum of the precision matrix is bounded.
Assumption 4.3
The precision matrix $\ \Omega^{\star}=(\Sigma^{\star})^{-1}$ has bounded operator $\ell_{1}$-norm and its eigenvalues are uniformly bounded away from 00 and $\infty$, i.e., there exists constants $\kappa\in(1,\infty)$ and $L\in(0,\infty)$ such that
$\Omega^{\star}\in\{\Omega\succeq 0\colon 1/\kappa\leq\lambda_{\mathrm{min}}(\Omega)\leq\lambda_{\mathrm{max}}(\Omega)\leq\kappa,\ \|\Omega\|_{\ell_{1}}\leq L\}.$
The boundedness assumption on the maximum absolute row sum of $\Omega^{\star}$ was also used by Yuan (2010). In our analysis it is only used to ensure that the residuals $\varepsilon_{j}$ have bounded sub-Gaussian norms, and it would not be necessary in the case of Gaussian data since then the tail probabilities can be controlled via the variances $\tau_{j}^{2}$ which are bounded due to the eigenvalue assumption.
Defining
$s=\mathop{\bf card}\{(j,k)\colon\Omega^{\star}_{jk}\neq 0,\ j>k\}$
as number of non-zero off-diagonal elements of the precision matrix $\Omega^{\star}$ (or, equivalently, the partial correlation matrix $Q^{\star}$), we can now state the main result of this section using the assumptions above.
Theorem 4.1
Under Assumptions 4.14.3 there exist positive constants $c$, $C_{1}$ and $C_{2}$ such that Algorithm 3.1 with $\lambda=c\sqrt{\log(p)/n}$ outputs an estimate $\widehat{Q}$ of the partial correlation matrix that satisfies
$\|\widehat{Q}-Q^{\star}\|_{\rm F}\leq C_{1}\sqrt{\frac{s\log p}{n}}$ (5)
and an estimate $\widehat{\Omega}$ of the precision matrix that satisfies
$\|\widehat{\Omega}-\Omega^{\star}\|_{\rm F}\leq C_{2}\sqrt{\frac{(s+p)\log p}{n}}$ (6)
with probability at least $1-6/p$.
The statistical rate of convergence (5) for the partial correlation matrix estimate is an improvement over the rates for the SPACE method of Peng et al. (2009) and the quasi-likelihood method of Khare et al. (2015). Although they are similarly $\sqrt{s}\lambda$, they require a choice of regularization parameter $\lambda$ that is asymptotically larger than ours, unless $s$ is bounded in which case they achieve a similar rate. The rate (6) for the precision matrix estimate matches that of the graphical lasso proved by Rothman et al. (2008). To the best of our knowledge, these are the best known Frobenius norm rates for partial correlation estimation and precision matrix estimation, respectively.
An inspection of the proof of Theorem 4.1 reveals that we can in fact obtain an explicit bound on the Frobenius error of the partial correlation matrix using quantities that, apart from $s$, can be computed in practice.
Corollary 4.2
Under the assumptions of Theorem 4.1, if $\lambda=c\sqrt{\log(p)/n}$ with $c$ sufficiently large, the estimate $\widehat{Q}$ produced by Algorithm 3.1 satisfies
$\|\widehat{Q}-Q^{\star}\|_{\rm F}\leq\frac{4}{\lambda_{\mathrm{min}}(\widehat{\Sigma})}\left(\frac{\max_{j}\hat{\tau}_{j}}{\min_{j}\hat{\tau}_{j}}\right)^{3}\sqrt{s}\lambda$
with probability at least $1-6/p$, where $\widehat{\Sigma}=(1/n)X^{\top}X$ is the sample covariance matrix.
5 A proximal splitting algorithm
To numerically compute the joint partial regression estimate, consider the equivalent problem to (3),
$\begin{array}[]{ll}\text{minimize}&\displaystyle f(\Omega)+g(\Omega)+h(\Omega)\end{array}$
where
$f(\Omega)=\sum_{j=1}^{p}\ell(X_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}\Omega_{\setminus j,j}),$
is the sum of the partial regression losses,
$g(\Omega)=\begin{cases}0&\text{if }\alpha I\preceq\Omega\preceq\beta I,\\ \infty&\text{otherwise}\end{cases}$
is the indicator function for the set $\mathcal{S}=\{\Omega\in\mbox{\bf S}^{p}\colon\alpha I\preceq\Omega\preceq\beta I\}$, and
$h(\Omega)=\sum_{j=1}^{p}\left(\lambda_{j}\hat{\tau}_{j}^{2}\|\Omega_{\setminus j,j}\|_{1}+\delta_{\{1/\hat{\tau}_{j}^{2}\}}(\Omega_{jj})\right),$
is the sum of the $\ell_{1}$-regularization terms and the indicator functions of the diagonal elements. Assuming $\ell$ is $L$-smooth for some $L\geq 0$, the general joint partial regression problem (3) can be solved via the PD3O algorithm (Yan, 2018). In this instance, the PD3O algorithm has iteration updates
$\Omega^{(k+1)}$ $=\mathbf{Prox}_{\gamma g}\left(\Omega^{(k)}-\gamma U^{(k)}-\gamma\nabla f(\Omega^{(k)})\right),$
$U^{(k+1)}$ $=\mathbf{Prox}_{\eta h^{*}}\left(U^{(k)}+\eta(2\Omega^{(k+1)}-\Omega^{(k)})+\gamma\eta(\nabla f(\Omega^{(k)})-\nabla f(\Omega^{(k+1)}))\right).$
The iterates $\Omega^{(k)}$ and $U^{(k)}$ are guaranteed to converge to a primal and a dual solution of the problem, respectively, if $\gamma$ and $\eta$ are chosen as positive step sizes satisfying $\gamma<2/L$ and $\gamma\eta\leq 1$.
Primal update.
The primal iterate update is a projection onto the set $\mathcal{S}$,
$\Omega^{(k+1)}=\Pi_{\mathcal{S}}\left(\Omega^{(k)}-\gamma U^{(k)}-\gamma\nabla f(\Omega^{(k)})\right).$
Importantly, note that the matrix which is being projected is not in general symmetric. The projection of a matrix $A\in{\mbox{\bf R}}^{n\times n}$ onto $\mathcal{S}$ is by definition given by
$\Pi_{\mathcal{S}}(A)=\operatornamewithlimits{argmin}_{\Omega\in\mathcal{S}}\|\Omega-A\|_{\rm F}^{2}.$
Due to symmetry, for any $\Omega\in\mathcal{S}$ the squared distance $\|\Omega-A\|_{\rm F}^{2}$ can be written as
$\sum_{j=1}^{p}(\Omega_{jj}-A_{jj})^{2}+2\sum_{j=1}^{p}\sum_{k<j}^{p}(\Omega_{jk}-(A_{jk}+A_{kj})/2)^{2}=\|\Omega-(A+A^{\top})/2\|_{\rm F}^{2}$ (7)
up to a term constant in $\Omega$. Let $Q\Lambda Q^{\top}=(A+A^{\top})/2$ be the eigenvalue decomposition of the symmetric part of $A$. The Frobenius norm is invariant under orthogonal transformations, so (7) can equivalently be written as
$\|Q^{\top}(\Omega-(A+A^{\top})/2)Q\|_{\rm F}^{2}=\|Q^{\top}\Omega Q-\Lambda\|_{\rm F}^{2}.$
Clearly, this expression is minimized by choosing $\Omega$ such that $Q^{\top}\Omega Q$ is diagonal, with the diagonal elements being the projection of the eigenvalues of ($A+A^{\top})/2$ onto the interval $[\alpha,\beta]$. Thus the projection $\Pi_{\mathcal{S}}(A)$ is given by projecting the symmetric part of $A$,
$\Pi_{\mathcal{S}}(A)$ $=Q\Lambda_{[\alpha,\beta]}Q^{\top}$
where $\Lambda_{[\alpha,\beta]}$ is a diagonal matrix with entries $(\Lambda_{[\alpha,\beta]})_{jj}=\min(\max(\Lambda_{jj},\alpha),\beta)$.
The gradient that appears in the primal step is given by
$\nabla f(\Omega)_{jj}$ $=0$
$\nabla f(\Omega)_{\setminus j,j}$ $=\hat{\tau}_{j}^{2}X_{\setminus j}^{\top}\nabla\ell(X_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}\Omega_{\setminus j,j}),$
and depends on the loss function we choose. If we take the quadratic loss as in (1), then for $i=1,\dots,n$,
$\nabla\ell(z)_{i}=\frac{z_{i}}{n}.$
For the Huber loss (2), we have instead
$\nabla\ell(z)_{i}=\frac{1}{n}\begin{cases}z_{i},&\text{if }|z_{i}|\leq\rho,\\ \rho\mathop{\bf sign}(z_{i}),&\text{otherwise}.\end{cases}$
For two points $\Omega$ and $\widetilde{\Omega}$ in ${\mbox{\bf R}}^{n\times n}$, with either loss function, $f$ satisfies
$\|\nabla f(\Omega)-\nabla f(\widetilde{\Omega})\|_{\rm F}$ $\leq\frac{1}{n}\left(\sum_{j=1}^{p}\|\hat{\tau}_{j}^{4}X_{\setminus j}^{\top}X_{\setminus j}(\Omega_{\setminus j,j}-\widetilde{\Omega}_{\setminus j,j})\|_{2}^{2}\right)^{1/2}$
$\leq\frac{1}{n}\max_{1\leq j\leq p}\hat{\tau}_{j}^{4}\lambda_{\text{max}}(X_{\setminus j}^{\top}X_{\setminus j})\|\Omega-\widetilde{\Omega}\|_{\rm F}.$
So $f$ is $L$-smooth with $L=\max_{1\leq j\leq p}\hat{\tau}_{j}^{4}\lambda_{\text{max}}(X_{\setminus j}^{\top}X_{\setminus j})/n$.
Dual update.
By the Moreau identity, we can write the dual iterate update with the proximal operator of $h/\eta$ instead of the Fenchel conjugate $\eta h^{*}$,
$V^{(k+1)}$ $=U^{(k)}+\eta(2\Omega^{(k+1)}-\Omega^{(k)})+\gamma\eta(\nabla f(\Omega^{(k)})-\nabla f(\Omega^{(k+1)})),$
$U^{(k+1)}$ $=V^{(k+1)}-\eta\mathbf{Prox}_{h/\eta}(V^{(k+1)}/\eta).$
The proximal operator of $h/\eta$ is given by
$\mathbf{Prox}_{h/\eta}(V)_{jj}$ $=1/\hat{\tau}_{j}^{2},$
$\mathbf{Prox}_{h/\eta}(V)_{kj}$ $=\mathop{\bf sign}(V_{kj})(V_{kj}-\hat{\tau}_{j}^{2}\lambda_{j}/\eta)_{+},$
for $j=1,\dots,p$ and $k\neq j$.
Termination criterion.
To stop the algorithm within a finite number of steps, we terminate when the iterates are approximately stationary. That is, given a tolerance $\epsilon^{\text{tol}}>0$, the termination criterion is
$\max(\|\Omega^{(k+1)}-\Omega^{(k)}\|_{\rm F},\|U^{(k+1)}-U^{(k)}\|_{\rm F})\leq\epsilon^{\text{tol}}.$
Computational complexity.
The computational cost of the projection in the primal update is that of an eigenvalue decomposition, which is $O(p^{3})$. In the case of the quadratic loss the gradient computation is $O(p^{3})$ if we cache the products $X_{\setminus j}^{\top}X_{j}$ and $X_{\setminus j}^{\top}X_{\setminus j}$. This means that apart from an initial matrix multiplication, the computational cost of the algorithm is independent of the number of samples $n$. For the Huber loss, however, the gradient computation is $O(np^{2})$, as we have to check the residual for each observation. The computational cost of the dual iterate update is $O(p^{2})$, meaning the total per iteration cost of the algorithm is $O(p^{3})$ when using the quadratic loss (matching the graphical lasso), and $O(p^{3}+np^{2})$ when using the Huber loss. Compared to the SPACE method, which has an iteration cost of $O(\min(p^{3},np^{2}))$, our algorithm is slightly more expensive for $n<p$. However, note that the SPACE method does not necessarily produce a negative semi-definite estimate, and that enforcing any eigenvalue constraints will necessarily incur a cost of $O(p^{3})$ due to the necessary eigenvalue decomposition.
6 Numerical experiments
We study the performance of the proposed method and compare it to existing estimation methods on synthetic data, and test the proposed method on real stock market data. In §6.1 we evaluate the performance of the estimators on synthetic data from a known distribution, meaning the true precision matrix is known. In §6.2 we provide a comparison of computation times in CPU seconds, and in §6.3 we estimate the partial correlation network of the returns of stocks on the Nasdaq Stock Market.
6.1 Synthetic data
We begin by evaluating the performance of the joint partial regression method on synthetic data for which the distribution and the true precision matrix are known. We compare the performance of the proposed method with the graphical lasso, as well as an oracle estimator which maximizes the Gaussian likelihood with the knowledge of the true sparsity pattern.
Data generation.
For different numbers of features $p$ we generate data sets of $n=500$ samples from a $p$-dimensional Gaussian distribution $\mathcal{N}(0,\Sigma)$. We generate the precision matrix $\Omega=\Sigma^{-1}$ according to three different models. In the first model we generate an Erdös–Rényi network with edge probability 0.05. That is, the adjacency matrix $A\in{\mbox{\bf R}}^{p\times p}$ is generated as
$A_{jk}=A_{kj}=\begin{cases}0&\text{with probability }0.95,\\ 1&\text{with probability }0.05.\\ \end{cases}$
In the second model we generate a simple path network, with the adjacency matrix $A$ having ones on the secondary diagonals. We denote this model by AR(1), since it corresponds to a first order autoregressive model. In the third model we generate a hub network as in Peng et al. (2009), where one node has a high degree of connectivity (connected to 20% of all other nodes), and the rest of the nodes have between one and three connections. Such a model is interesting in the context of gene regulatory networks, where a single gene may regulate a large number of other genes.
In all three models we generate the precision matrix by randomly flipping the signs of the adjacency matrix and setting the diagonal elements of matrix to $\Omega_{jj}=1+\sum_{k\neq j}|A_{jk}|$ to guarantee positive definiteness via strict diagonal dominance.
Regularization parameter selection.
We select the regularization parameters for the graphical lasso and joint partial regression via 5-fold cross-validation. For joint partial regression we use different regularization parameters $\lambda_{j}$ for each feature, and select them separately via cross-validation in the initial regression step.
Results.
We compare the average Frobenius and operator $\ell_{2}$-norm error of the three methods over 100 generated data sets. The average Frobenius errors are shown in Figure 1, and the average operator $\ell_{2}$-norm are shown in Figure 2. The results show that the joint partial regression method outperforms the graphical lasso almost uniformly across all models and problem sizes, with Figure 2(c) even showing joint partial regression as closer to the oracle estimator than it is to the graphical lasso.
(a) Erdös–Rényi model.
(b) $\text{AR}(1)$ model.
(c) Hub network model.
Figure 1: Average Frobenius error versus number of features with ±$\pm 2\times\text{SE}$ bands.
(a) Erdös–Rényi model.
(b) $\text{AR}(1)$ model.
(c) Hub network model.
Figure 2: Average operator $\ell_{2}$-error versus number of features with ±$\pm 2\times\text{SE}$ bands.
6.2 Timing comparisons
We compare the average timings of our implementation of joint partial regression and an implementation of the graphical lasso over $10$ datasets of $n=500$ Gaussian samples with different numbers of features $p$. The timings are shown in Table 1. Both methods are run with $\lambda=10^{-2}$, which gives them a similar level of sparsity.
Joint partial regression is implemented in a Python package written in Rust, using the proximal algorithm derived in §5 along with FISTA (Beck and Teboulle, 2009) for the initial regression step. The graphical lasso is implemented in the Python package Scikit-learn (Pedregosa et al., 2011), and uses the algorithm proposed by Friedman et al. (2008) coupled with a coordinate descent algorithm written in Cython.
Note that the termination criteria of the methods are disanalogous, since the graphical lasso uses a dual residual while joint partial regression uses the difference of iterates. Hence we compare timings both when the methods are run until the tolerance $\epsilon^{\text{tol}}=10^{-3}$ and until the 100th iteration is reached, respectively. All tests were performed on an Intel Core i9 3.2 GHz processor.
Table 1: Average timings (CPU seconds) with ±×$\pm\times\mathrm{SE}$ for joint partial regression and graphical lasso on ten random datasets of $p$ features and $n=500$ observations.
(a) Termination after tolerance $\epsilon^{\text{tol}}=10^{-3}$ is reached.
$p$ Joint partial regression Graphical lasso
100 $\mathbf{0.11\pm 0.024}$ $1.01\pm 0.227$
200 $\mathbf{0.64\pm 0.038}$ $5.91\pm 0.706$
400 $\mathbf{5.09\pm 0.091}$ $29.81\pm 3.737$
(b) Termination after 100 iterations.
$p$ Joint partial regression Graphical lasso
100 $\mathbf{0.31\pm 0.022}$ $1.67\pm 0.278$
200 $\mathbf{1.33\pm 0.095}$ $7.29\pm 0.977$
400 $\mathbf{6.98\pm 0.499}$ $42.76\pm 1.916$
6.3 Stock market data
To illustrate the performance of the proposed method on real data, we estimate the partial correlation network of stocks on the Nasdaq Stock Market. The data consists of the daily returns of 399 out of the 500 highest market capitalization stocks listed on stock exchanges in the United States. The data spans from the 11th of February 2013 to the 7th of February 2018, and contains 1258 observations. The network, estimated using the proposed method, is shown in Figure 3. Here the nodes represent the stocks and the color indicates the sector to which the stock belongs. We can clearly see structure relating to sector belonging in the network. The four largest connected components of the network are overwhelmingly comprised of stocks from the finance (magenta), utilities (red), real estate (green), and energy (orange) sectors, respectively.
Figure 3: Stock market network as estimated by the proposed method.
7 Discussion
We propose a method for estimating high-dimensional sparse precision matrices and partial correlation matrices based on jointly regressing each feature on the remaining features. The method is shown to match the statistical rate of convergence for precision matrix estimation of the graphical lasso, and to improve upon the rate for the partial correlation matrix estimation of the SPACE method. We also derive an efficient algorithm for solving the joint partial regression problem based on the PD3O algorithm. The method was shown to perform well on both synthetic and real data, outperforming the graphical lasso in terms of both estimation error and computation time.
Appendix A Proof of Theoretical Results
A.1 Technical lemmas
Lemma A.1
Set $\theta_{j}^{\star}=-\tau_{j}^{2}\Omega_{\setminus j,j}^{\star}$ and let
$\hat{\theta}_{j}=\operatornamewithlimits{argmin}_{\theta_{j}}\left\{\frac{1}{2n}\|X_{j}-X_{\setminus j}\theta_{j}\|_{2}^{2}+\lambda\|\theta_{j}\|_{1}\right\}$
be the lasso estimate of $\theta_{j}^{\star}$. Suppose that Assumptions 4.24.1 holds, the design matrix satisfies
$\gamma_{\ell}\|\nu\|_{2}^{2}\leq\frac{1}{n}\|X\nu\|_{2}^{2}\leq\gamma_{u}\|\nu\|_{2}^{2}\quad\text{for all }\nu\in{\mbox{\bf R}}^{p},$
for some $\gamma_{\ell},\gamma_{u}\in(0,\infty)$, and $\lambda=c\sqrt{\log(p)/n}$ for some sufficiently large constant $c>0$. Then, with probability at least $1-4/p$, the estimates $\hat{\tau}_{j}^{2}=(1/n)\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}$ of the residual variances satisfy
$\max_{j}|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\lesssim\sqrt{\frac{\log p}{n}},$ (8)
and
$\tau_{\ell}\leq\hat{\tau}_{j}\leq\tau_{u}\quad\text{for all }j=1,\dots,p,$ (9)
for some constants $\tau_{\ell},\tau_{u}\in(0,\infty)$, and the regularization parameter satisfies
$\lambda\geq\max\bigg{(}$ $\max_{j}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\gamma_{u}\kappa|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\right)$ (10)
$\max_{j}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\kappa\frac{\gamma_{u}}{\tau_{\ell}}\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|\right)\bigg{)}.$
Proof  Recall that $X_{j}=X_{\setminus j}\theta_{j}^{\star}+\epsilon_{j}$, where $\epsilon_{j}$ is an $n$-dimensional random variable with zero mean. From Bickel et al. (2009), we have that if $\lambda\geq(2/n)\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}$, then
$\|\hat{\theta}_{j}-\theta_{j}^{\star}\|_{2}\leq\frac{3}{\gamma_{\ell}}\sqrt{s_{j}}\lambda,\quad\|\hat{\theta}_{j}-\theta_{j}^{\star}\|_{1}\leq\frac{12}{\gamma_{\ell}}s_{j}\lambda,$ (11)
and
$\frac{1}{n}\|X_{\setminus j}(\hat{\theta}-\theta^{\star})\|_{2}^{2}\leq\frac{9}{\gamma_{\ell}}s_{j}\lambda^{2}.$ (12)
where $s_{j}=\mathop{\bf card}(\theta_{j})$. Note that $X_{k}$ for $k\neq j$ and $\epsilon_{j}$ are uncorrelated, so ${\mathbb{E}}(X_{k}^{\top}\epsilon_{j})=0$. Moreover, $\epsilon_{j}=X_{j}-X_{\setminus j}\theta_{j}$,
$\|\epsilon_{ij}\|_{\psi_{2}}\leq\|X_{ij}\|_{\psi_{2}}+\sum_{k\neq j}\|X_{ik}\theta_{jk}^{\star}\|_{\psi_{2}}\lesssim(1+\|\theta_{j}^{\star}\|_{1})K\leq(1+\kappa\|\Omega^{\star}\|_{\ell_{1}})K$
by the triangle inequality of the sub-Gaussian norm. So, due to Assumption 4.3, for all $i$ and $j$, $\epsilon_{ij}$ is sub-Gaussian with norm bounded by $(1+\kappa L)K$, and thus for all $k\neq j$, $X_{ik}\epsilon_{ij}$ is sub-exponential, meaning there exists a constant $C$ such that for every $t>0$,
${\mathbb{P}}\left(\frac{1}{n}\left|\sum_{i=1}^{n}X_{ik}\epsilon_{ij}\right|\geq t\right)\leq 2\exp(-t/C).$
The sub-exponential norm of a random variable $Y$ is defined as
$\|Y\|_{\psi_{1}}=\inf\left\{t>0\colon{\mathbb{E}}(\exp(Y/t))\leq 2\right\}.$
Then, by (Vershynin, 2018, Lemma 2.7.7), $\|X_{ik}\epsilon_{ij}\|_{\psi_{1}}\leq\|X_{ik}\|_{\psi_{2}}\|\epsilon_{ij}\|_{\psi_{2}}\leq K^{\prime}$, where $K^{\prime}=(1+\kappa L)K^{2}$. By Bernstein’s inequality (Vershynin, 2018, Corollary 2.8.3),
${\mathbb{P}}\left(\frac{2}{n}|X_{k}^{\top}\epsilon_{j}|\geq\lambda\right)\leq 2\exp\left(-c_{0}\left(\frac{\lambda}{K^{\prime}}\right)^{2}n\right),$
whenever $\log(p)/n$ is sufficiently small, and where $c_{0}$ is a universal constant. Hence,
${\mathbb{P}}\left(\lambda\leq\max_{j}\frac{2}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\right)$ $\leq\sum_{j=1}^{p}{\mathbb{P}}\left(\frac{2}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\geq\lambda\right)$
$\leq\sum_{j=1}^{p}\sum_{k\neq j}{\mathbb{P}}\left(\frac{2}{n}|X_{k}^{\top}\epsilon_{j}|\geq\lambda\right)$
$\leq 2p(p-1)\exp\left(-c_{0}\left(\frac{\lambda}{K^{\prime}}\right)^{2}n\right)$
$=2\exp\left(-c_{0}\left(\frac{\lambda}{K^{\prime}}\right)^{2}n+\log p+\log(p-1)\right).$
Then, choosing $\lambda=c\sqrt{\log(p)/n}$ with $c>0$ sufficiently large ensures that $\lambda\geq(2/n)\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}$ holds for all $j$ with probability at least $1-2/p$. So the inequalities (11) and (12) hold with same probability for this choice of $\lambda$. Turning to the residual variances, we have
$\hat{\tau}_{j}^{2}-\tau_{j}^{2}$ $=\frac{1}{n}\|X_{\setminus j}(\hat{\theta}_{j}-\theta_{j})\|_{2}^{2}+\frac{2}{n}\langle\epsilon_{j},X_{\setminus j}(\hat{\theta}_{j}-\theta_{j})\rangle+\left(\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right)$
$\leq\frac{9}{\gamma_{\ell}}s_{j}\lambda^{2}+\frac{2}{n}\langle\epsilon_{j},X_{\setminus j}(\hat{\theta}_{j}-\theta_{j})\rangle+\left(\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right).$
Then, noting that
$\frac{2}{n}|\langle\epsilon_{j},X_{\setminus j}(\hat{\theta}_{j}-\theta_{j})\rangle|\leq\frac{2}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\hat{\theta}_{j}-\theta_{j}\|_{1}\leq\frac{12}{\gamma_{\ell}}s_{j}\lambda^{2},$
where the first inequality is due to Hölder and the second is due to (11), we get the bound
$|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\leq\frac{21}{\gamma_{\ell}}s_{j}\lambda^{2}+\left|\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right|=c^{2}\frac{21}{\gamma_{\ell}}\frac{s_{j}\log p}{n}+\left|\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right|.$ (13)
The random variable $(1/n)\|\epsilon_{j}\|_{2}^{2}$ is sub-exponential with mean $\tau_{j}^{2}$ and $\|\epsilon_{ij}\|_{\psi_{1}}\leq(1+\kappa L)^{2}K^{2}$, so another application of Bernstein’s inequality yields
${\mathbb{P}}\left(\left|\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right|\geq\frac{(1+\kappa L)^{2}K^{2}}{\sqrt{c_{0}/2}}\sqrt{\frac{\log p}{n}}\right)\leq\sum_{j=1}^{p}2\exp\left(-2\log p\right),$
and so,
${\mathbb{P}}\left(\left|\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right|\geq\frac{(1+\kappa L)^{2}K^{2}}{\sqrt{c_{0}/2}}\sqrt{\frac{\log p}{n}}\text{ for some }j\right)$ $\leq\sum_{j=1}^{p}2\exp\left(-2\log p\right)$
$=2p\exp(-2\log p)=2/p.$
Due to Assumption 4.1, $s_{j}\log(p)/n\leq M\sqrt{\log(p)/n}$, so we have that $|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\lesssim\sqrt{\log(p)/n}$ for all $j$ with probability at least $1-4/p$, proving (8). Due to this bound, $\hat{\tau}_{j}^{2}=\tau_{j}^{2}+O(\sqrt{\log(p)/n})$, so since $\lambda_{\mathrm{min}}(\Omega^{\star})\leq 1/\tau_{j}^{2}\leq\lambda_{\mathrm{max}}(\Omega^{\star})$, there exists $\tau_{\ell},\tau_{u}\in(0,\infty)$ such that $\tau_{\ell}\leq\hat{\tau}_{j}\leq\tau_{u}$ for all $j\in\{1,\dots,p\}$, whenever $\log(p)/n$ is small, proving (9).
Since $\lambda=c\sqrt{\log(p)/n}\geq(2/n)\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}$ for all $j$, we have that the regularization parameter satisfies
$\lambda\geq\max_{j}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\gamma_{u}\kappa|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\right)$
with the same probability, if $c$ is sufficiently large. Likewise, since
$\sqrt{\frac{\log p}{n}}\gtrsim|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|=|\hat{\tau}_{j}+\tau_{j}||\hat{\tau}_{j}-\tau_{j}|\geq(\tau_{u}+\tau_{j})|\hat{\tau}_{j}-\tau_{j}|,$
it holds that
$\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|$ $=\max_{k}\left|\frac{\hat{\tau}_{j}\tau_{k}-\tau_{j}\hat{\tau}_{k}}{\hat{\tau}_{k}\tau_{k}}\right|$
$=\max_{k}\left|\frac{\left(\tau_{j}+O(\sqrt{\log(p)/n})\right)\tau_{k}-\tau_{j}\left(\tau_{k}+O(\sqrt{\log(p)/n})\right)}{\hat{\tau}_{k}\tau_{k}}\right|$
$\lesssim\sqrt{\frac{\log p}{n}}.$
Thus, if $c$ is sufficiently large, then the inequality
$\lambda\geq\max_{j}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\kappa\frac{\gamma_{u}}{\tau_{\ell}}\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|\right)$
also holds with the same probability.
Lemma A.2
For any vector $\nu\in{\mbox{\bf R}}^{p}$, the data matrix satisfies
$\lambda_{\mathrm{min}}(\Sigma)\left(1-\sqrt{\frac{p+2\log p}{n}}\right)^{2}\|\nu\|_{2}^{2}\leq\frac{1}{n}\|X\nu\|_{2}^{2}\leq\lambda_{\mathrm{max}}(\Sigma)\left(1+\sqrt{\frac{p+2\log p}{n}}\right)^{2}\|\nu\|_{2}^{2},$
with probability at least $1-2/p$.
Proof  Let $\Sigma=Q\Lambda Q^{\top}$ be the eigenvalue decomposition of the covariance matrix. Then $X=Q\Lambda^{1/2}A$ where $A$ is a matrix of $n$ i.i.d. rows of mean-zero, sub-Gaussian and isotropic vectors in ${\mbox{\bf R}}^{p}$. Hence, $X^{\top}X=\Lambda^{1/2}Y^{\top}Y\Lambda^{1/2}$. Without loss of generality, take any unit vector $\nu\in{\mbox{\bf R}}^{p}$. Then
$\frac{1}{n}\|X\nu\|_{2}^{2}$ $=\frac{1}{n}\|Q\Lambda^{1/2}Y\nu\|_{2}^{2}=\frac{1}{n}\|\Lambda^{1/2}Y\nu\|_{2}^{2}$
$\geq\lambda_{\mathrm{min}}(\Lambda)\lambda_{\mathrm{min}}((1/n)Y^{\top}Y)=\lambda_{\mathrm{min}}(\Sigma)\lambda_{\mathrm{min}}((1/n)Y^{\top}Y).$
The upper bound
$\frac{1}{n}\|X\nu\|_{2}^{2}\leq\lambda_{\mathrm{max}}(\Sigma)\lambda_{\mathrm{max}}((1/n)Y^{\top}Y)$
follows similarly. A straightforward application of (Vershynin, 2011, Corollary 5.35) then gives the result.
A.2 Proof of Theorem 4.1
Proof  Note that under Assumption 4.1, Lemma A.2 tells us that there exists $\gamma_{\ell},\gamma_{u}\in(0,\infty)$ such that $\gamma_{\ell}\|\nu\|_{2}^{2}\leq(1/n)\|X\nu\|_{2}^{2}\leq\gamma_{u}\|\nu\|_{2}^{2}$ for all $\nu\in{\mbox{\bf R}}^{p}$ with probability $1-2/p$. Thus, by Lemma A.1, inequalities (8) and (10) hold with probability at least $1-6/p$. Throughout, assume that this event occurs.
Define the matrix $\Omega^{\prime}$ by
$\Omega_{jj}^{\prime}=1/\hat{\tau}_{j}^{2},\quad\Omega_{jk}^{\prime}=\Omega_{jk}^{\star},\quad j\neq k.$ (14)
Then we have that
$\lambda_{\text{min}}(\Omega^{\prime})$ $=\lambda_{\text{min}}(\Omega-\mathop{\bf diag}(\Omega-\widehat{\Omega}))$ (15)
$\geq\lambda_{\text{min}}(\Omega)-\max_{j=1,\dots,p}|1/\hat{\tau}_{j}^{2}-1/\tau_{j}^{2}|.$
By Assumption 4.3 and Lemma A.1, $\lambda_{\text{min}}(\Omega^{\prime})\geq 0$ whenever $\log(p)/n$ is sufficiently small, meaning $\Omega^{\prime}$ is feasible in the joint partial regression problem (1). To alleviate the notation, let $\omega^{\star}_{j}=\Omega_{\setminus j,j}^{\star}$, and set $G$ as the objective function of (1) except with shifted variables, i.e.,
$G(\nu_{1},\dots,\nu_{p})=\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}(\omega^{\star}_{j}+\nu_{j})\|_{2}^{2}+\lambda\hat{\tau}^{2}_{j}\|\omega^{\star}_{j}+\nu_{j}\|_{1}\right).$
By construction, we have the inequality $G(\hat{\nu}_{1},\dots,\hat{\nu}_{p})\leq G(0,\dots,0)$, where $\hat{\nu}_{j}=\widehat{\Omega}_{\setminus j,j}-\omega^{\star}_{j}$. Recall that
$X_{j}=-\tau_{j}^{2}X_{\setminus j}\omega^{\star}_{j}+\epsilon_{j},\quad\epsilon_{j}\sim\mathcal{N}(0,\tau_{j}^{2}I_{n}),$
Expanding some squares, we have that
$G(0,\dots,0)$ $=\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}\omega^{\star}_{j}\|_{2}^{2}+\lambda\hat{\tau}_{j}^{2}\|\omega^{\star}_{j}\|_{1}\right)$ (16)
$=\sum_{j=1}^{p}\left(\frac{1}{2n}\|\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j}\|_{2}^{2}+\lambda\hat{\tau}_{j}^{2}\|\omega^{\star}_{j}\|_{1}\right)$
and,
$G(\hat{\nu}_{1},\dots,\hat{\nu}_{p})$ $=\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}(\omega^{\star}_{j}+\hat{\nu}_{j})\|_{2}^{2}+\lambda\hat{\tau}_{j}^{2}\|\omega^{\star}_{j}+\hat{\nu}_{j}\|_{1}\right)$
$=\sum_{j=1}^{p}\left(\frac{1}{2n}\|\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}\hat{\nu}_{j}\|_{2}^{2}+\lambda\hat{\tau}_{j}^{2}\|\omega^{\star}_{j}+\hat{\nu}_{j}\|_{1}\right)$
$=\sum_{j=1}^{p}\bigg{(}\frac{1}{2n}\|\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j}\|_{2}^{2}+\frac{\hat{\tau}_{j}^{2}}{n}\langle\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j},X_{\setminus j}\hat{\nu}_{j}\rangle$
$+\frac{1}{2n}\|\hat{\tau}_{j}^{2}X_{\setminus j}\hat{\nu}_{j}\|_{2}^{2}+\lambda\hat{\tau}_{j}^{2}\|\omega^{\star}_{j}+\hat{\nu}_{j}\|_{1}\bigg{)}$
Rearranging some terms and using $(1/n)\|X\nu\|_{2}^{2}\geq\gamma_{\ell}\|\nu\|_{2}^{2}$, we obtain the inequality
$\sum_{j=1}^{p}\frac{\gamma_{\ell}\hat{\tau}_{j}^{2}}{2}\|\hat{\nu}_{j}\|_{2}^{2}\leq\sum_{j=1}^{p}\left(\lambda\hat{\tau}_{j}^{2}(\|\omega^{\star}_{j}\|_{1}-\|\omega^{\star}_{j}+\hat{\nu}_{j}\|_{1})-\frac{\hat{\tau}_{j}^{2}}{n}\langle\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j},X_{\setminus j}\hat{\nu}_{j}\rangle\right).$
Using Hölder’s inequality and $\lambda_{\text{max}}((1/n)X^{\top}X)\leq\gamma_{u}$ and $\lambda_{\text{min}}(\Sigma)\geq 1/\kappa$, we have that
$\left|\frac{1}{n}\langle\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j},X_{\setminus j}\hat{\nu}_{j}\rangle\right|$ $=\left|\frac{1}{n}\langle\epsilon_{j},X_{\setminus j}\hat{\nu}_{j}\rangle+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})\frac{1}{n}\langle X_{\setminus j}\omega^{\star}_{j},X_{\setminus j}\hat{\nu}_{j}\rangle\right|$
$\leq\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\hat{\nu}_{j}\|_{1}+|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\frac{1}{n}\|X^{\top}X\omega^{\star}_{j}\|_{2}\|\hat{\nu}_{j}\|_{2}$
$\leq\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\hat{\nu}_{j}\|_{1}+\gamma_{u}\kappa|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\|\hat{\nu}_{j}\|_{1}$
$=\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\gamma_{u}\kappa|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\right)\|\hat{\nu}_{j}\|_{1}.$
Let $S_{j}$ be the indices of the true non-zero elements of $\omega^{\star}_{j}$, and $S_{j}^{c}$ be the indices of the zero elements. Then we have
$\|\omega^{\star}_{j}+\hat{\nu}_{j}\|_{1}=\|(\omega^{\star}_{j}+\hat{\nu}_{j})_{S_{j}}\|_{1}+\|(\hat{\nu}_{j})_{S_{j}^{c}}\|_{1}\geq\|\omega^{\star}_{j}\|_{1}-\|(\hat{\nu}_{j})_{S_{j}}\|_{1}$ (17)
where we subtracted $\|(\hat{\nu}_{j})_{S_{j}^{c}}\|_{1}$ and used the reverse triangle inequality. Using (17) and (10), we have that
$\sum_{j=1}^{p}\frac{\gamma_{\ell}}{2}\|\hat{\tau}_{j}^{2}\hat{\nu}_{j}\|_{2}^{2}\leq\sum_{j=1}^{p}\left(\frac{3}{2}\lambda\hat{\tau}_{j}^{2}\|(\hat{\nu}_{j})_{S_{j}}\|_{1}\right)$ (18)
Thus, we have that
$\left(\frac{\gamma_{\ell}}{2}\min_{j}\hat{\tau}_{j}^{4}\right)\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F}^{2}\leq\left(\frac{3}{2}\max_{j}\hat{\tau}_{j}^{2}\right)\lambda\|(\widehat{\Omega}-\Omega^{\prime})_{S}\|_{1}\leq\left(\frac{3}{2}\max_{j}\hat{\tau}_{j}^{2}\right)\lambda\sqrt{s}\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F},$ (19)
where $S$ are the indices of the true non-zero entries. By some rearranging, this gives the bound
$\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F}\leq\left(\frac{3}{\gamma_{\ell}}\frac{\max_{j}\hat{\tau}^{2}_{j}}{\min_{j}\hat{\tau}^{4}_{j}}\right)\sqrt{s}\lambda.$ (20)
By the inequality (8), $\hat{\tau}_{j}^{2}=\tau_{j}^{2}+O(\sqrt{\log(p)/n})$, so since $\lambda_{\mathrm{min}}(\Omega^{\star})\leq 1/\tau_{j}^{2}\leq\lambda_{\mathrm{max}}(\Omega^{\star})$, there exists $\tau_{\ell},\tau_{u}\in(0,\infty)$ such that $\tau_{\ell}\leq\hat{\tau}_{j}\leq\tau_{u}$ for all $j\in\{1,\dots,p\}$, whenever $\log(p)/n$ is small. Hence, the term $\max_{j}\hat{\tau}^{2}_{j}/\min_{j}\hat{\tau}^{4}_{j}$ is bounded, so (20) implies that $\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F}\lesssim\sqrt{s\log(p)/n}$. Moreover,
$|1/\hat{\tau}_{j}^{2}-1/\tau_{j}^{2}|\leq\frac{1}{\hat{\tau}_{j}^{2}\tau_{j}^{2}}|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\leq\frac{\lambda_{\mathrm{max}}(\Omega^{\star})}{\tau_{\ell}^{2}}|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\lesssim\sqrt{\frac{\log p}{n}},$ (21)
Hence,
$\|\widehat{\Omega}-\Omega^{\star}\|_{\rm F}^{2}=\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F}^{2}+\sum_{j=1}^{p}(1/\hat{\tau}_{j}^{2}-1/\tau_{j}^{2})^{2}\lesssim\frac{(s+p)\log p}{n},$
which is the stated result for the precision matrix.
Turning to the partial correlation matrix, we define the function $H$ by
$H(\nu_{1},\dots,\nu_{p})=\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}-X_{\setminus j}\widehat{R}_{j}(q^{\star}_{j}+\nu_{j})\|_{2}^{2}+\lambda\|\widehat{R}_{j}(q^{\star}_{j}+\nu_{j})\|_{1}\right),$
where $q_{j}^{\star}=Q_{\setminus j,j}^{\star}$ and $\widehat{R}_{j}=\mathop{\bf diag}(\hat{\tau}_{j}/\hat{\tau}_{1},\dots,\hat{\tau}_{j}/\hat{\tau}_{j-1},\hat{\tau}_{j}/\hat{\tau}_{j+1},\dots,\hat{\tau}_{j}/\hat{\tau}_{p})$. Like before, we have that
$H(0,\dots,0)$ $=\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}-X_{\setminus j}\widehat{R}_{j}q^{\star}_{j}\|_{2}^{2}+\lambda\hat{\tau}_{j}^{2}\|\widehat{R}_{j}q^{\star}_{j}\|_{1}\right)$ (22)
$=\sum_{j=1}^{p}\left(\frac{1}{2n}\|\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j}\|_{2}^{2}+\lambda\|\widehat{R}_{j}q^{\star}_{j}\|_{1}\right)$
where $R_{j}=\mathop{\bf diag}(\tau_{j}/\tau_{1},\dots,\tau_{j}/\tau_{j-1},\tau_{j}/\tau_{j+1},\dots\tau_{j}/\tau_{p})$. Likewise,
$H(\tilde{\nu}_{1},\dots,\tilde{\nu}_{p})$ $=\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}-\hat{X}_{\setminus j}\widehat{R}_{j}(q^{\star}_{j}+\tilde{\nu}_{j})\|_{2}^{2}+\lambda\|\widehat{R}_{j}(q^{\star}_{j}+\tilde{\nu}_{j})\|_{1}\right)$
$=\sum_{j=1}^{p}\left(\frac{1}{2n}\|\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j}-X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\|_{2}^{2}+\lambda\|\widehat{R}_{j}(q^{\star}_{j}+\tilde{\nu}_{j})\|_{1}\right)$
$=\sum_{j=1}^{p}\bigg{(}\frac{1}{2n}\|\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j}\|_{2}^{2}-\frac{1}{n}\langle\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j},X_{\setminus j}\widehat{R}_{j}\hat{\nu}_{j}\rangle$
$\quad+\frac{1}{2n}\|X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\|_{2}^{2}+\lambda\|\widehat{R}_{j}(q^{\star}_{j}+\tilde{\nu}_{j})\|_{1}\bigg{)}$
The true partial correlation matrix $Q^{\star}$ is always feasible in the joint partial regression problem, so for $\tilde{\nu}_{j}=\widehat{Q}_{\setminus j,j}-q_{j}^{\star}$, we have that $H(\tilde{\nu}_{1},\dots,\tilde{\nu}_{p})\leq H(0,\dots,0)$. By rearranging terms, we have that
$\sum_{j=1}^{p}\frac{\gamma_{\ell}}{2}\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{2}^{2}\leq\sum_{j=1}^{p}\Big{(}$ $\lambda(\|\widehat{R}_{j}q^{\star}_{j}\|_{1}-\|\widehat{R}_{j}(q^{\star}_{j}+\tilde{\nu}_{j})\|_{1})$
$+\frac{1}{n}\langle\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j},X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\rangle\Big{)}.$
Due to $\widehat{R}_{j}$ being diagonal, it holds that $\lambda_{\mathrm{min}}(\widehat{R}_{j})=\min_{k\neq j}\hat{\tau}_{j}/\hat{\tau}_{k}\geq\tau_{\ell}/\tau_{u}$ and $\|\widehat{R}_{j}\|_{\ell_{1}}=\lambda_{\mathrm{max}}(\widehat{R}_{j})=\max_{k\neq j}\hat{\tau}_{j}/\hat{\tau}_{k}\leq\tau_{u}/\tau_{\ell}$. Thus, we get the inequality
$\frac{\gamma_{\ell}(\tau_{\ell}/\tau_{u})^{2}}{2}\sum_{j=1}^{p}\|\tilde{\nu}_{j}\|_{2}^{2}\leq\sum_{j=1}^{p}\Big{(}$ $\lambda(\tau_{u}/\tau_{\ell})(\|q^{\star}_{j}\|_{1}-\|q^{\star}_{j}+\tilde{\nu}_{j}\|_{1})$ (23)
$+\frac{1}{n}\langle\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j},X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\rangle\Big{)}.$
Once again using Hölder’s inequality, we have that the inner product terms satisfy
$\left|\frac{1}{n}\langle\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j},X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\rangle\right|$ $=\left|\frac{1}{n}\langle\epsilon_{j},X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\rangle-\frac{1}{n}\langle X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j},X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\rangle\right|$
$\leq\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{1}+\frac{1}{n}\|X^{\top}X(\widehat{R}_{j}-R_{j})q^{\star}_{j}\|_{2}\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{2}$
$\leq\frac{\tau_{u}}{\tau_{\ell}}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\tilde{\nu}_{j}\|_{1}+\gamma_{u}\lambda_{\mathrm{max}}(\widehat{R}_{j}-R_{j})\|q^{\star}_{j}\|_{2}\|\tilde{\nu}_{j}\|_{2}\right)$
$\leq\frac{\tau_{u}}{\tau_{\ell}}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\tilde{\nu}_{j}\|_{1}+\kappa\frac{\gamma_{u}}{\tau_{\ell}}\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|\|\tilde{\nu}_{j}\|_{1}\right)$
$=\frac{\tau_{u}}{\tau_{\ell}}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\kappa\frac{\gamma_{u}}{\tau_{\ell}}\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|\right)\|\tilde{\nu}_{j}\|_{1}.$
Thus, using (10), the inequality (23) implies that
$\sum_{j=1}^{p}\frac{\gamma_{\ell}}{2}\left(\frac{\tau_{\ell}}{\tau_{u}}\right)^{2}\|\tilde{\nu}_{j}\|_{2}^{2}\leq\frac{\tau_{u}}{\tau_{\ell}}\sum_{j=1}^{p}\lambda(\|q^{\star}_{j}\|_{1}-\|q^{\star}_{j}+\tilde{\nu}_{j}\|_{1}+\|\tilde{\nu}_{j}\|_{1})$
Adapting the inequality (17), we have that
$\sum_{j=1}^{p}\frac{\gamma_{\ell}}{2}\left(\frac{\tau_{\ell}}{\tau_{u}}\right)^{2}\|\tilde{\nu}_{j}\|_{2}^{2}\leq 2\frac{\tau_{u}}{\tau_{\ell}}\sum_{j=1}^{p}\lambda\|(\tilde{\nu}_{j})_{S_{j}}\|_{1}.$
By rearranging terms and using the fact that $\sum_{j}\|\tilde{\nu}_{j}\|_{2}^{2}=\|\widehat{Q}-Q^{\star}\|_{\rm F}^{2}$ and $\sum_{j}\|\tilde{\nu}_{j}\|_{1}=\|\widehat{Q}-Q^{\star}\|_{1}$, we get that
$\|\widehat{Q}-Q^{\star}\|_{\rm F}^{2}\leq\frac{4}{\gamma_{\ell}}\left(\frac{\tau_{u}}{\tau_{\ell}}\right)^{3}\lambda\|(\widehat{Q}-Q^{\star})_{S}\|_{1}\leq\frac{4}{\gamma_{\ell}}\frac{\tau_{\ell}}{\tau_{u}}\lambda\sqrt{s}\|\widehat{Q}-Q^{\star}\|_{\rm F}.$
Inserting $\lambda=c\sqrt{\log(p)/n}$ and dividing by the Frobenius error, we arrive at the bound
$\|\widehat{Q}-Q^{\star}\|_{\rm F}\leq\frac{4c}{\gamma_{\ell}}\left(\frac{\tau_{u}}{\tau_{\ell}}\right)^{3}\sqrt{\frac{s\log p}{n}},$
completing the proof.