### v2

# Inverse Covariance and Partial Correlation Matrix Estimation via Joint Partial Regression

Abstract
We present a method for estimating sparse high-dimensional inverse covariance and partial correlation matrices, which exploits the connection between the inverse covariance matrix and linear regression. The method is a two-stage estimation method wherein each individual feature is regressed on all other features while positive semi-definiteness is enforced simultaneously. We derive non-asymptotic estimation rates for both inverse covariance and partial correlation matrix estimation. An efficient proximal splitting algorithm for numerically computing the estimate is also dervied. The effectiveness of the proposed method is demonstrated on both synthetic and real-world data.
1 Introduction
Two important and closely related problems in statistical learning are the problems of estimating a partial correlation network and the inverse covariance matrix, also known as the precision matrix, from data. Partial correlation networks, which generalize the Gaussian graphical model, are used to model the relationships between variables when controlling for other variables. Partial correlation networks are used in a plethora of applications, e.g., the analysis of gene expression data (de la Fuente et al.,, 2004), quantitative psychology (Epskamp and Fried,, 2018), and the analysis of financial markets (Kenett et al.,, 2015). The precision matrix, from which we can obtain the partial correlation network, is also of interest in its own right, as it appears in linear discriminant analysis (Hastie et al.,, 2009) and in Markowitz portfolio selection (Markowitz,, 1952). However, due to the high-dimensionality of the problem, estimating a precision or partial correlation matrix is often challenging as the number of parameters is on the order of the squared number of features. For this reason, classical methods such as using the inverse of the sample covariance matrix, are known to perform poorly whenever the number of observation is not extremely large. Additionally, classical methods produce estimates which are almost surely dense. This makes regularization crucial, since in many applications we typically only have a moderate number of observations, and since we typically seek a sparse estimate which gives rise to a more parsimonious and interpretable network model.
Related work.
Gaussian graphical models, and by extension sparse precision matrix estimation, have been studied extensively in the literature, with the covariance selection problem originally being due to Dempster, (1972). Today there are many methods for estimating the precision matrix, which can be divided into three main categories: methods based upon maximum likelihood estimation, methods based upon approximate inversion, and methods based upon what we shall refer to as partial regression. The partial regression approach exploits a connection between the precision matrix and linear regression, and is given its name due to its close relation to partial correlations.
Perhaps the most popular method of estimation is $\ell_{1}$-regularized Gaussian maximum likelihood estimation, commonly known as the graphical lasso. This method was introduced independently by Yuan and Lin, (2007) and d’Aspremont et al., (2008) (having slight variations), where the former work also provided a consistency result in the high-dimensional setting. There has additionally been much work on the problem of numerically solving the graphical lasso problem, such as the work of Friedman et al., (2008), who also coined the name. A more recent addition to methods in this vein is given by Bertsimas et al., (2020), which directly solves the cardinality constrained maximum likelihood problem to optimality using mixed-integer programming.
Cai et al., (2011) provide a method of the second kind, clime, based on minimizing the $\ell_{1}$-norm of the estimate under the constraint that the estimate is sufficiently close to an inverse of the sample covariance matrix. Robust extensions of this method are proposed by Zhao and Liu, (2014) and Wang and Gu, (2017), focusing on heavy-tailed and contaminated data, respectively. The work of Cai et al., (2016) provide an adaptive variant, aclime. They also derive the minimax optimal rate in the $\ell_{2}$-operator norm, and show that which their methods achieves it.
The present work is closely related to the works of Meinshausen and Bühlmann, (2006), van de Geer et al., (2014), and Yuan, (2010), who opt for estimation via partial regression. Meinshausen and Bühlmann, specifically consider the sparsity pattern estimation problem in a high-dimensional setting, and use the lasso (Tibshirani,, 1996) to estimate the neighborhood of each node in the graphical model. Their work provides consistency results of their proposed method, and has been one of the most impactful works on the subject. The work of van de Geer et al., continues in the same vein and estimate the precision matrix with the purpose of constructing confidence regions for high-dimensional models. Yuan, provides a method that can be boiled down to solving a sequence of linear programs, opting for the Dantzig selector (Candès and Tao,, 2007) in lieu of the lasso, and projecting the estimate onto the space of symmetric matrices with respect to the operator $\ell_{1}$-norm. The method is proven to achieve the minimax optimal rate in the operator $\ell_{1}$-norm over certain parameter spaces.
There are additionally methods that combine partial regression with pseudo-likelihood to estimate the precision matrix. The first method in this vein is the symmetric lasso procedure due to Friedman et al., (2010). Perhaps the most recent contribution is given by Behdin et al., (2023), who pursue this approach with discrete optimization, as they opt for direct penalization of the cardinality, rather than its usual convex surrogate. Behdin et al., also give a non-asymptotic error rate for their estimator.
Less work has been done on the problem of estimating high-dimensional partial correlation matrices, despite their frequent use in data analysis. Peng et al., (2009) provide an alternating minimization method dubbed space for estimating the partial correlation matrix via partial regression. Khare et al., (2015) also propose concord, a pseudo-likelihood partial regression method which is similar to the symmetric lasso procedure. Peng et al., and Khare et al., both provide asymptotic analysis for their respective methods, achieving identical error in the limit up to a constant.
Contributions.
The main contributions of the present work are three-fold, and are as follows:
  • We propose a method for simultaneously estimating a sparse precision and partial correlation matrix via linear regression, which we call joint partial regression. For sub-Gaussian data, we provide non-asymptotic estimation error rates, which match the best known rate for precision matrix estimation and improve upon the best known asymptotic rate for partial correlation matrix estimation. To the best of our knowledge, our non-asymptotic result on partial correlation estimation is the first of its kind.
  • We provide an efficient proximal algorithm for numerically fitting the proposed method which is well-suited for large-scale data sets. An implementation of the algorithm is also given in the form of a Python package written in Rust.
  • We conduct numerical studies on synthetic data, showing that the proposed method outperforms the graphical lasso almost uniformly over several precison matrix models and problem sizes. We also apply the method to stock market data, showing that it is able to produce interpretable results on real data.
Notation.
We write $\mbox{\bf S}^{p}$ for the set of real symmetric $p\times p$ matrices, and for $\Omega\in\mbox{\bf S}^{p}$ we write $\Omega\succeq 0$ if $\Omega$ is positive semi-definite. If $\Omega$ and $\Omega^{\prime}$ are both in $\mbox{\bf S}^{p}$ we write $\Omega\succeq\Omega^{\prime}$ if $\Omega-\Omega^{\prime}\succeq 0$. The Frobenius norm is defined as $\|\Omega\|_{\rm F}=\sqrt{\mathop{\bf Tr}(\Omega^{2})}$. The cardinality function $\mathop{\bf card}$ returns the number of elements if given a set argument, and the number of non-zero entries if given a vector argument. The proximal operator of a closed, convex and proper function $f$ is denoted $\mathbf{Prox}_{f}$.
2 The precision matrix and partial correlation
Likely the most common use of the precision matrix is in the context of Gaussian graphical models, where the connection between the precision matrix and the conditional independence structure of jointly Gaussian random variables is exploited. The basis of this connection is in fact a connection that the precision matrix also has to linear regression. Suppose $Z$ is a square-integrable mean-zero random vector, taking values in ${\mbox{\bf R}}^{p}$, with non-singular covariance $\Sigma$ and inverse $\Omega=\Sigma^{-1}$. Then the best linear unbiased predictor of a component $Z_{j}$ given all other components
$Z_{\setminus j}=(Z_{1},\dots,Z_{j-1},Z_{j+1},\dots,Z_{p})$
is $Z_{\setminus j}^{\top}\theta_{j}$, where
$\theta_{j}=\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j}.$
Here $\Sigma_{\setminus j,\setminus j}$ is the sub-matrix of $\Sigma$ obtained by removing the $j$-th row and column, and $\Sigma_{\setminus j,j}$ is the $j$-th column of $\Sigma$ with the $j$-th element removed. Likewise, the residual $\varepsilon_{j}=Z_{j}-Z_{\setminus j}^{\top}\theta_{j}$ has mean zero and variance $\tau_{j}^{2}=\Sigma_{jj}-\Sigma_{j,\setminus j}\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j}$. Using Schur complements, we have that
$\Omega_{jj}$ $=(\Sigma_{jj}-\Sigma_{j,\setminus j}\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j})^{-1},$
$\Omega_{\setminus j,j}$ $=-(\Sigma_{jj}-\Sigma_{j,\setminus j}\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j})^{-1}\Sigma_{\setminus j,\setminus j}^{-1}\Sigma_{\setminus j,j},$
and thus
$\Omega_{jj}$ $=1/\tau_{j}^{2},$ (2.1)
$\Omega_{\setminus j,j}$ $=-\theta_{j}/\tau_{j}^{2}.$
In the Gaussian case we find that $\varepsilon_{j}$ is independent of $Z_{\setminus j}$. This then implies that $Z_{j}$ and $Z_{k}$ are conditionally independent, given $Z_{\setminus\{j,k\}}$ if and only if $\theta_{jk}=0$, i.e., $\Omega_{jk}=0$. More generally, the partial correlation coefficient between $Z_{j}$ and $Z_{k}$ given $Z_{\setminus\{j,k\}}$ is given by the negative correlation between $\varepsilon_{j}$ and $\varepsilon_{k}$,
$\rho_{jk\mid\setminus\{j,k\}}=-\mathop{\bf Corr}(\varepsilon_{j},\varepsilon_{k})=-\frac{{\mathbb{E}}(\varepsilon_{j}\varepsilon_{k})}{\sqrt{{\mathbb{E}}(\varepsilon_{j}^{2}){\mathbb{E}}(\varepsilon_{k}^{2})}},$
which can be written in terms of the precision matrix and the linear regression coefficients as
$\rho_{jk\mid\setminus\{j,k\}}=-\frac{\Omega_{jk}}{\sqrt{\Omega_{jj}\Omega_{kk}}}=\frac{\tau_{k}}{\tau_{j}}\theta_{jk},$
respectively. Thus we define the partial correlation matrix $Q$ as
$Q=-T\Omega T,$
where $T=\mathop{\bf diag}(\tau_{1},\dots,\tau_{p})$. (Note that some works (Schäfer and Strimmer,, 2004; Marrelec et al.,, 2006) define the partial correlation matrix as $-T\Omega T+2I$, so that the diagonal elements are ones.)
3 Proposed method
In light of the connection between the precision matrix and linear regression described in §2, a way to estimate the precision matrix is to form estimates of the linear regression coefficients $\hat{\theta}_{j}$ and the residual variances $\hat{\tau}_{j}^{2}$ for each $j$ and then compute an estimate of the precision matrix $\widehat{\Omega}$ via
$\widehat{\Omega}_{jj}$ $=1/\hat{\tau}_{j}^{2},$
$\widehat{\Omega}_{\setminus j,j}$ $=-\hat{\theta}_{j}/\hat{\tau}_{j}^{2},$
and an estimate of the partial correlation matrix $\widehat{Q}$ via
$\widehat{Q}=-\widehat{T}\widehat{\Omega}\widehat{T},$
where $\widehat{T}=\mathop{\bf diag}(\hat{\tau}_{1},\dots,\hat{\tau}_{p})$. An obvious approach, since we are interested in sparsity, is to use the lasso. Suppose we have access to a data matrix $X\in{\mbox{\bf R}}^{n\times p}$, where each row contains a sample of the random vector $Z$ and each column corresponds to a feature of the data. We can then estimate the linear regression coefficients by solving the lasso problem
$\hat{\theta}_{j}=\operatornamewithlimits{argmin}_{\theta\in\mathbf{R}^{p-1}}\left\{\frac{1}{2n}\|X_{j}-X_{\setminus j}\theta\|_{2}^{2}+\lambda\|\theta\|_{1}\right\}$
and subsequently estimate the residual variances by $\hat{\tau}_{j}^{2}=(1/n)\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}$. However, this approach does not result in a symmetric estimate of the precision matrix, let alone a positive semi-definite one. If we are solely interested in the unweighted partial correlation network, we could choose to include the edge $(j,k)$ only if $\widehat{\Omega}_{jk}$ and $\widehat{\Omega}_{kj}$ are both non-zero or, alternatively, if either is non-zero. If we are interested in the partial correlations themselves we could choose to symmetrize the estimate by averaging the estimates of $\widehat{\Omega}_{jk}$ and $\widehat{\Omega}_{kj}$, or by projection onto the cone of positive semi-definite matrices in some norm. We will however take a different approach, in which we enforce positive semi-definiteness while simultaneously estimating the linear regression coefficients, given initial estimates of the residual variances.
3.1 Joint partial regression
We propose the following convex program that regresses the features at the same time as ensuring positive semi-definiteness,
$\begin{array}[]{ll}\text{minimize}&\displaystyle\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}-X_{\setminus j}\theta_{j}\|_{2}^{2}+\lambda\|\theta_{j}\|_{1}\right)\\ \text{subject to}&\Omega_{jj}=1/\hat{\tau}_{j}^{2},\quad\Omega_{\setminus j,j}=-\theta_{j}/\hat{\tau}_{j}^{2},\quad j=1,\dots,p,\\ &\Omega\succeq 0,\quad Q=-\widehat{T}\Omega\widehat{T}.\end{array}$ (3.1)
Here $\Omega\in{\mathbf{R}}^{p\times p}$ and $\theta_{j}\in{\mathbf{R}}^{p-1}$, $j=1,2,\dots,p$, are the optimization variables, and the data are the matrix $X$, the estimates $\hat{\tau}_{j}^{2}$ of the residual variances, and the regularization parameter $\lambda\geq 0$. Note that the equality $Q=-\widehat{T}\Omega\widehat{T}$ is not a real constraint as $Q$ does not appear in the objective function or other constraints, but simply expresses that $Q$ is a by-product of the optimization problem in the stated way. We call the solutions $\widehat{\Omega}$ and $\widehat{Q}$ the joint partial regression estimates of the precision matrix and the partial correlation matrix, respectively. In order to obtain the initial estimates of the residual variances we can use the aforementioned lasso estimates of the linear regression coefficients to compute the sample variances $\hat{\tau}_{j}^{2}=(1/n)\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}$. We summarize the proposed procedure in Algorithm 3.1.
Note moreover that applying the derivation in §2 to the sample covariance matrix reveals that Algorithm 3.1 exactly retrieves the inverse sample covariance matrix when $\lambda=0$, if it exists.
Input: Data matrix $X\in{\mbox{\bf R}}^{n\times p}$, penalty parameter $\lambda$.
for $j=1,\dots,p$ do
lasso regression
$\hat{\theta}_{j}=\operatornamewithlimits{argmin}_{\theta\in{\mathbf{R}}^{p-1}}\left\{\frac{1}{2n}\|X_{j}-X_{\setminus j}\theta\|_{2}^{2}+\lambda\|\theta\|_{1}\right\},$
and compute the estimate $\hat{\tau}_{j}^{2}=(1/n)\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}$ of the residual variance.
end
Solve the convex program (3.1) with the estimated residual variances $\hat{\tau}_{j}^{2}$ and regularization parameter $\lambda$ to obtain the estimates $\widehat{\Omega}$ and $\widehat{Q}$ of the precision matrix and partial correlation matrix, respectively.
Algorithm 3.1 Joint partial regression
3.2 Proximal splitting algorithm
Consider a generalization to problem (3.1), where we allow for any convex and $L$-smooth loss $\ell$, a different regularization parameter $\lambda_{j}$ for each feature $j$, and eigenvalue limits $\alpha$ and $\beta$:
$\begin{array}[]{ll}\text{minimize}&\displaystyle f(\Omega)+g(\Omega)\\ \text{subject to}&\alpha I\preceq\Omega\preceq\beta I\end{array}$ (3.2)
where
$f(\Omega)=\sum_{j=1}^{p}\ell(X_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}\Omega_{\setminus j,j})$
is the sum of partial regression losses, and
$g(\Omega)=\sum_{j=1}^{p}\left(\lambda_{j}\hat{\tau}_{j}^{2}\|\Omega_{\setminus j,j}\|_{1}+\delta_{\{1/\hat{\tau}_{j}^{2}\}}(\Omega_{jj})\right),$
is the sum of $\ell_{1}$-regularization terms and indicator functions of the diagonal elements. The generalized joint partial regression problem (3.2) can be solved via the pd3o algorithm (Yan,, 2018). In this instance, the pd3o algorithm has iteration updates
$\Omega^{(k+1)}$ $=\Pi_{\mathcal{S}}\left(\Omega^{(k)}-\gamma U^{(k)}-\gamma\nabla f(\Omega^{(k)})\right),$
$U^{(k+1)}$ $=\mathbf{Prox}_{\eta g^{*}}\left(U^{(k)}+\eta(2\Omega^{(k+1)}-\Omega^{(k)})+\gamma\eta(\nabla f(\Omega^{(k)})-\nabla f(\Omega^{(k+1)}))\right).$
where $\Pi_{\mathcal{S}}$ is the Euclidean projection onto the set $\mathcal{S}=\{\Omega\in\mbox{\bf S}^{p}\colon\alpha I\preceq\Omega\preceq\beta I\}$, and $g^{*}$ is the Fenchel conjugate of $g$. The iterates $\Omega^{(k)}$ and $U^{(k)}$ are guaranteed to converge to a primal and a dual solution of the problem, respectively, if $\gamma$ and $\eta$ are chosen as positive step sizes satisfying $\gamma<2/L$ and $\gamma\eta\leq 1$. We provide more details on the computation of the primal and dual updates in Appendix B.
Termination criterion.
To stop the algorithm within a finite number of steps, we terminate when the iterates are approximately stationary. That is, given a tolerance $\epsilon^{\text{tol}}>0$, the termination criterion is
$\max\left(\|\Omega^{(k+1)}-\Omega^{(k)}\|_{\rm F},\|U^{(k+1)}-U^{(k)}\|_{\rm F}\right)\leq\epsilon^{\text{tol}}.$
Computational complexity.
The computational cost of the projection in the primal update is that of an eigenvalue decomposition, which is $O(p^{3})$. In the case of the quadratic loss the gradient computation is $O(p^{3})$ if we initially cache the matrix products, and $O(p^{2}n)$ otherwise. The computational cost of the dual update is $O(p^{2})$, meaning the total per iteration cost of the algorithm is $O(p^{3})$ when using the quadratic loss, matching the graphical lasso’s complexity.
Practical considerations.
The initial lasso regressions in Algorithm 3.1 may give us several beneficial by-products. First, forming a matrix via (2.1) makes for an appropriate initialization $\Omega^{(0)}$. Second, we may use cross-validation or information criterions, such as the Akaike or Bayesian information criterions, to select the regularization parameters before solving the optimization problem (3.2). In this way, we only need to run $p$ lasso parameter selection procedures to select a parameter for each feature, which would be prohibitively expensive if instead done on the final estimate level. Moreover, another benefit of the partial regression approach is the we can easily swap out the quadratic loss for another, e.g. the Huber loss, in the case of contaminated data.
4 Theoretical results
In this section we provide theoretical guarantees for the estimation error of the proposed method that hold with high probability under given assumptions. Each data point, a random vector in ${\mbox{\bf R}}^{p}$, is assumed to have a true, unknown, covariance matrix that we denote by $\Sigma^{\star}$. The corresponding true precision and partial correlation matrix is $\Omega^{\star}$ and $Q^{\star}$, respectively. Our first assumption concerns how the dimensionality of the problem grows. This first assumption specifies how fast $p$ may grow with $n$, and also bounds the growth of the degree of the true partial correlation network.
Assumption 4.1.
The dimensionality is such that $p/n\leq 1-\delta$ for some $\delta\in(0,1)$, and the degree $d=\max_{j}\sum_{k\neq j}\mathbf{1}(\Omega_{jk}^{\star}\neq 0)$ of the partial correlation network is such that $d\sqrt{\log(p)/n}\leq M$ for some constant $M>0$.
Note that the dependence on $n$ and $p$ is in many cases not made explicit; e.g., in the above assumption both $\Omega^{\star}$ and $d$ tacitly depend on $n$ and $p$. We remark moreover that the term $d\sqrt{\log(p)/n}$ that appears in the assumption is interesting in its own right. It is the minimax error rate in the operator $\ell_{q}$-norm for $1\leq q\leq\infty$ as shown by Cai et al., (2016), and it is the error rate of the graphical lasso in the operator $\ell_{2}$-norm (the spectral norm), as shown by Ravikumar et al., (2011). Noting that the spectral norm is smaller than the Frobenius norm, we argue that the boundedness assumption is comparatively weak.
Our second assumption is that the data is sub-Gaussian. A real-valued random variable $Y$ is said to be sub-Gaussian if there exists a constant $C$ such that the tail probabilities satisfy
${\mathbb{P}}(|Y|\geq t)\leq 2\exp(-t^{2}/C^{2}),\quad\text{ for all }t\geq 0,$
and a multivariate random variable $Z$ is said to be sub-Gaussian if every linear combination $\langle z,Z\rangle$ with $z\in{\mbox{\bf R}}^{p}$ is sub-Gaussian. We define the sub-Gaussian norm $\|\cdot\|_{\psi_{2}}$ for real-valued random variables by
$\|Y\|_{\psi_{2}}=\inf\{t>0\colon{\mathbb{E}}(\exp(Y^{2}/t^{2}))\leq 2\},$
and for multivariate random variables by $\|Z\|_{\psi_{2}}=\sup_{\|z\|_{2}=1}\|\langle z,Z\rangle\|_{\psi_{2}}$.
Assumption 4.2.
The rows $X_{i}$ of the data matrix $X\in{\mbox{\bf R}}^{n\times p}$ are $n$ i.i.d. samples drawn from a sub-Gaussian distribution with covariance matrix $\Sigma^{\star}$ and sub-Gaussian norm $\|X_{i}\|_{\psi_{2}}\leq K$ for some constant $K>0$.
It is clear from Assumption 4.2 that each element $X_{ij}$ is sub-Gaussian with $\|X_{ij}\|_{\psi_{2}}\leq K$. We remark however that the assumption that the row-wise norms $\|X_{i}\|_{\psi_{2}}$ stay bounded as $p$ grows does not imply that the element-wise norms $\|X_{ij}\|_{\psi_{2}}$ need to go to zero. This is because the multivariate sub-Gaussian norm measures, loosely speaking, the size of the largest eigenvalue of the covariance matrix (cf. Vershynin,, 2018, for the Gaussian case). Thus, as long as this eigenvalue stays bounded, which is part of our next assumption, there is generelly no such tacit assumption on the $X_{ij}$.
Finally we assume that the eigenvalues of the true covariance matrix do not degenerate to zero or infinity.
Assumption 4.3.
The eigenvalues of the precision matrix $\Omega^{\star}=(\Sigma^{\star})^{-1}$ are uniformly bounded away from 00 and $\infty$, i.e., there exists a constant $\kappa\in(1,\infty)$ such that $1/\kappa\leq\lambda_{\mathrm{min}}(\Omega^{\star})\leq\lambda_{\mathrm{max}}(\Omega^{\star})\leq\kappa.$
Defining the size $s$ of the partial correlation network as the number of non-zero off-diagonal elements of the precision matrix $\Omega^{\star}$,
$s=\mathop{\bf card}\{(j,k)\colon\Omega^{\star}_{jk}\neq 0,\ j>k\},$
we can now state the main result of this section using the assumptions above.
Theorem 4.1.
Under Assumptions 4.1, 4.2 and 4.3, there exist positive constants $c$, $C_{1}$ and $C_{2}$ such that Algorithm 3.1 with $\lambda=c\sqrt{\log(p)/n}$ outputs an estimate $\widehat{Q}$ of the partial correlation matrix that satisfies
$\|\widehat{Q}-Q^{\star}\|_{\rm F}\leq C_{1}\sqrt{\frac{s\log p}{n}}$ (4.1)
and an estimate $\widehat{\Omega}$ of the precision matrix that satisfies
$\|\widehat{\Omega}-\Omega^{\star}\|_{\rm F}\leq C_{2}\sqrt{\frac{(s+p)\log p}{n}}$ (4.2)
with probability at least $1-6/p$.
The statistical rate of convergence (4.1) for the partial correlation matrix estimate is an improvement over the asymptotic rates of space (Peng et al.,, 2009) and concord (Khare et al.,, 2015). Although they are similarly $\sqrt{s}\lambda$, they require a choice of regularization parameter $\lambda$ that is asymptotically larger than ours, unless $s$ is bounded in which case they achieve a similar rate. The rate (4.2) for the precision matrix estimate matches that of the graphical lasso proved by Rothman et al., (2008). To the best of our knowledge, these are the best known Frobenius norm rates for partial correlation and precision matrix estimation, respectively. We however acknowledge that the theoretical analysis considers the case $p<n\ll p^{2}$, and not the case when there are fewer samples than features which other works have.
An inspection of the proof of Theorem 4.1 reveals that we can in fact obtain an explicit bound on the Frobenius error of the partial correlation matrix using quantities that, apart from the size $s$, can be computed in practice.
Corollary 4.2.
Under the assumptions of Theorem 4.1, if $\lambda=c\sqrt{\log(p)/n}$ with $c$ sufficiently large, the estimate $\widehat{Q}$ produced by Algorithm 3.1 satisfies
$\|\widehat{Q}-Q^{\star}\|_{\rm F}\leq\frac{4}{\lambda_{\mathrm{min}}(\widehat{\Sigma})}\left(\frac{\max_{j}\hat{\tau}_{j}}{\min_{j}\hat{\tau}_{j}}\right)^{3}\sqrt{s}\lambda$
with probability at least $1-6/p$, where $\widehat{\Sigma}=(1/n)X^{\top}X$ is the sample covariance matrix.
5 Numerical experiments
We study the performance of the proposed method and compare it to existing estimation methods on synthetic data, and test the proposed method on real stock market data. In §5.1 we evaluate the performance of the estimators on synthetic data from a known distribution. In §5.2 we provide a comparison of computation times in cpu seconds, and in §5.3 we estimate the partial correlation network of the returns of stocks on the Nasdaq Stock Market.
5.1 Synthetic data
We begin by evaluating the performance of the joint partial regression method on synthetic data for which the distribution and the true precision matrix are known. We compare the performance of the proposed method with the graphical lasso, as well as an oracle estimator which maximizes the Gaussian likelihood with the knowledge of the true sparsity pattern.
Data generation.
For different numbers of features $p$ we generate data sets of $n=500$ samples from a $p$-dimensional Gaussian distribution $\mathcal{N}(0,\Sigma)$. We generate the precision matrix $\Omega=\Sigma^{-1}$ according to three different models. In the first model we generate an Erdös–Rényi network with edge probability 0.05. That is, the adjacency matrix $A\in{\mbox{\bf R}}^{p\times p}$ is generated as
$A_{jk}=A_{kj}=\begin{cases}0&\text{with probability }0.95,\\ 1&\text{with probability }0.05.\\ \end{cases}$
In the second model we generate a simple path network, with the adjacency matrix $A$ having ones on the secondary diagonals. We denote this model by AR(1), since it corresponds to a first order autoregressive model. In the third model we generate a hub network as in Peng et al., (2009), where one node has a high degree of connectivity (connected to 20% of all other nodes), and the rest of the nodes have between one and three connections. Such a model is interesting in the context of gene regulatory networks, where a single gene may regulate a large number of other genes.
In all three models we generate the precision matrix by randomly flipping the signs of the adjacency matrix and setting the diagonal elements of matrix to $\Omega_{jj}=1+\sum_{k\neq j}|A_{jk}|$ to guarantee positive definiteness via strict diagonal dominance.
Regularization parameter selection.
We select the regularization parameters for the graphical lasso and joint partial regression via 5-fold cross-validation. For joint partial regression we use different regularization parameters $\lambda_{j}$ for each feature, and select them separately via cross-validation in the initial regression step, as described in §3.2.
Results.
We compare the average Frobenius and operator $\ell_{2}$-norm error of the three methods over 100 generated data sets. The average Frobenius errors are shown in Figure 1, and the average operator $\ell_{2}$-norm errors are shown in Figure 2. The results show that the joint partial regression method outperforms the graphical lasso almost uniformly across network models and problem sizes, with Figure 2(c) even showing joint partial regression as closer to the oracle estimator than it is to the graphical lasso.
(a) Erdös–Rényi model.
(b) $\text{AR}(1)$ model.
(c) Hub network model.
Figure 1: Average Frobenius error versus number of features with ±$\pm 2\times\text{SE}$ bands.
(a) Erdös–Rényi model.
(b) $\text{AR}(1)$ model.
(c) Hub network model.
Figure 2: Average operator $\ell_{2}$-error versus number of features with ±$\pm 2\times\text{SE}$ bands.
5.2 Timing comparisons
We compare the average timings of our implementation of joint partial regression and an implementation of the graphical lasso over $10$ datasets of $n=500$ Gaussian samples with different numbers of features $p$. The timings are shown in Table 1. Both methods are run with $\lambda=10^{-2}$, which gives them a similar level of sparsity.
Joint partial regression is implemented in a Python package written in Rust, using the proximal algorithm derived in §3.2 along with fista (Beck and Teboulle,, 2009) for the initial regression step. The graphical lasso is implemented in the Python package Scikit-learn (Pedregosa et al.,, 2011), and uses the algorithm proposed by Friedman et al., (2008) coupled with a coordinate descent algorithm written in Cython.
Note that the termination criteria of the methods are disanalogous, since the graphical lasso uses a dual residual while joint partial regression uses the difference of iterates. Hence we compare timings both when the methods are run until the tolerance $\epsilon^{\text{tol}}=10^{-3}$ and until the 100th iteration is reached. All tests were performed on an Intel Core i9 3.2 GHz processor.
Table 1: Average timings (cpu seconds) with $\pm$±SE for joint partial regression and graphical lasso on ten random datasets of $p$ features and $n=500$ observations.
(a) Termination after tolerance is reached.
$p$ Joint partial regression Graphical lasso
100 0.11 $\pm$ 0.024 1.01 $\pm$ 0.227
200 0.64 $\pm$ 0.038 5.91 $\pm$ 0.706
400 5.09 $\pm$ 0.091 29.81 $\pm$ 3.737
(b) Termination after 100 iterations.
Joint partial regression Graphical lasso
0.31 $\pm$ 0.022 1.67 $\pm$ 0.278
1.33 $\pm$ 0.095 7.29 $\pm$ 0.977
6.98 $\pm$ 0.499 42.76 $\pm$ 1.916
5.3 Stock market data
To illustrate the performance of the proposed method on real data, we estimate the partial correlation network of stocks on the Nasdaq Stock Market. The data (Nugent,, 2018) consists of daily prices of the 500 highest market capitalization stocks listed on stock exchanges in the United States. The data spans from the 11th of February 2013 to the 7th of February 2018, and contains 1258 observations. We calculate the daily returns, and after removing stocks with missing data and those without Nasdaq sector data, we are left with 399 assets. The network, estimated using the proposed method, is shown in Figure 3. Here the nodes represent the stocks and the color indicates the sector to which the stock belongs. We can clearly see structure relating to sector belonging in the network. The four largest connected components of the network are overwhelmingly comprised of stocks from the finance (magenta), utilities (red), real estate (green), and energy (orange) sectors, respectively.
Figure 3: Stock market network as estimated by the proposed method.
6 Discussion
We propose a method for estimating high-dimensional sparse precision and partial correlation matrices based on jointly regressing each feature on the remaining features. Our theoretical analysis of partial correlation estimation provides what is to the best of our knowledge the first non-asymptotic rate. The method is shown to match the rate of the graphical lasso for precision matrix estimation, and has an improved asymptotic error compared to space and concord. We also derive an efficient algorithm for solving the joint partial regression problem based on the pd3o algorithm. The method was shown to perform well on both synthetic and real data, producing interpretable results and outperforming the graphical lasso in terms of estimation error.
Appendix A Proof of Theoretical Results
Define the ${\mbox{\bf R}}^{n}$-valued random vector $\epsilon_{j}$ by the relation $X_{j}=X_{\setminus j}\theta_{j}^{\star}+\epsilon_{j}$, and note that its elements are independent zero mean random variables.
A.1 Technical lemmas
Lemma A.1.
Set $\theta_{j}^{\star}=-\tau_{j}^{2}\Omega_{\setminus j,j}^{\star}$ and let
$\hat{\theta}_{j}=\operatornamewithlimits{argmin}_{\theta_{j}}\left\{\frac{1}{2n}\|X_{j}-X_{\setminus j}\theta_{j}\|_{2}^{2}+\lambda\|\theta_{j}\|_{1}\right\}$
be the lasso estimate of $\theta_{j}^{\star}$. Suppose that Assumptions 4.1, 4.2 and 4.3 hold, the data matrix satisfies
$\gamma_{\ell}\|\nu\|_{2}^{2}\leq\frac{1}{n}\|X\nu\|_{2}^{2}\leq\gamma_{u}\|\nu\|_{2}^{2}\quad\text{for all }\nu\in{\mbox{\bf R}}^{p},$
for some $\gamma_{\ell},\gamma_{u}\in(0,\infty)$, and $\lambda=c\sqrt{(\log p)/n}$ for some sufficiently large constant $c>0$.
Then, with probability at least $1-4/p$, the estimates $\hat{\tau}_{j}^{2}=(1/n)\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}$ of the residual variances satisfy
$\max_{j}|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\lesssim\sqrt{\frac{\log p}{n}}$ (A.1)
and
$\tau_{\ell}\leq\hat{\tau}_{j}\leq\tau_{u}\quad\text{for all }j=1,\dots,p,$ (A.2)
for some constants $\tau_{\ell},\tau_{u}\in(0,\infty)$, and the regularization parameter satisfies
$\lambda\geq\max_{1\leq j\leq p}$ $\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\gamma_{u}\kappa|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\right)$ (A.3)
$\vee\max_{1\leq j\leq p}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\kappa\frac{\gamma_{u}}{\tau_{\ell}}\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|\right).$
Proof.
From Hastie et al., (2015) (see also Bickel et al.,, 2009, for a fuller background) we have that if $\lambda\geq(2/n)\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}$, then
$\|\hat{\theta}_{j}-\theta_{j}^{\star}\|_{2}\leq\frac{3}{\gamma_{\ell}}\sqrt{s_{j}}\lambda,\quad\|\hat{\theta}_{j}-\theta_{j}^{\star}\|_{1}\leq\frac{12}{\gamma_{\ell}}s_{j}\lambda,$ (A.4)
and
$\frac{1}{n}\|X_{\setminus j}(\hat{\theta}-\theta^{\star})\|_{2}^{2}\leq\frac{9}{\gamma_{\ell}}s_{j}\lambda^{2}.$ (A.5)
where $s_{j}=\mathop{\bf card}(\theta_{j})$. The first and third of these inequalities are Equations (11.14b) and (11.25b) respectively in Hastie et al., (2015), and the second inequality follows from the first one and the bound $\|\hat{\theta}-\theta^{\star}\|_{1}\leq 4s_{j}^{1/2}\|\hat{\theta}-\theta^{\star}\|_{2}$. This latter bound in turns follows from Hastie et al.,, the very end of the proof of its Theorem 1.1, and its Lemma 11 (in their notation): $\|\widehat{\nu}\|_{1}=\|\widehat{\nu}_{S}\|_{1}+\|\widehat{\nu}_{S^{c}}\|_{1}\leq\|\widehat{\nu}_{S}\|_{1}+3\|\widehat{\nu}_{S}\|_{1}=4\|\widehat{\nu}_{S}\|_{1}\leq 4k^{1/2}\|\widehat{\nu}\|_{2}$, with $\hat{\nu}$ and $k$ corresponding to our $\hat{\theta}-\theta^{\star}$ and $s_{j}$ respectively.
Now note that $\epsilon_{j}$ and $X_{k}$ are uncorrelated for $k\neq j$, so ${\mathbb{E}}(X_{k}^{\top}\epsilon_{j})=0$. Moreover, since $\epsilon_{j}=X_{j}-X_{\setminus j}\theta_{j}$ we have
$\|\epsilon_{ij}\|_{\psi_{2}}=\|X_{ij}-X_{i,\setminus j}\theta^{\star}_{j}\|_{\psi_{2}}=\tau_{j}^{2}\|\langle X_{i},\Omega_{j}^{\star}\rangle\|_{\psi_{2}}\leq\tau_{j}^{2}\|\Omega_{j}^{\star}\|_{2}K\leq\tau_{j}^{2}\lambda_{\max}(\Omega^{\star})K\leq\kappa^{2}K$
by Assumptions 4.2 and 4.3. Note also that we used that $\tau_{j}^{2}=1/\Omega_{jj}$ are bounded due to the eigenvalue assumption. Specifically,
$1/\kappa\leq\lambda_{\min}(\Omega^{\star})\leq\Omega_{jj}^{\star}=1/\tau_{j}^{2}$ (A.6)
for all $j$ and hence $\tau_{j}^{2}\leq\kappa$. This implies (Vershynin,, 2018, Lemma 2.7.7) that $X_{ik}\epsilon_{ij}$ is sub-exponential for all $k\neq j$. As a sum of such random variables is also sub-exponential, there exists a constant $C$ such that for every $t>0$,
${\mathbb{P}}\left(\frac{1}{n}\left|\sum_{i=1}^{n}X_{ik}\epsilon_{ij}\right|\geq t\right)\leq 2\exp(-t/C).$
The sub-exponential norm of a random variable $Y$ is defined as
$\|Y\|_{\psi_{1}}=\inf\left\{t>0\colon{\mathbb{E}}(\exp(Y/t))\leq 2\right\}.$
By Vershynin, (2018, Lemma 2.7.7), $\|X_{ik}\epsilon_{ij}\|_{\psi_{1}}\leq\|X_{ik}\|_{\psi_{2}}\|\epsilon_{ij}\|_{\psi_{2}}\leq\kappa^{2}K^{2}$. By Bernstein’s inequality (Vershynin,, 2018, Corollary 2.8.3),
${\mathbb{P}}\left(\frac{2}{n}|X_{k}^{\top}\epsilon_{j}|\geq\lambda\right)\leq 2\exp\left(-c_{0}\left(\frac{\lambda}{\kappa^{2}K^{2}}\right)^{2}n\right),$
whenever $(\log p)/n$ is sufficiently small and where $c_{0}$ is a universal constant. Hence
${\mathbb{P}}\left(\lambda\leq\max_{j}\frac{2}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\right)$ $\leq\sum_{j=1}^{p}{\mathbb{P}}\left(\frac{2}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\geq\lambda\right)$
$\leq\sum_{j=1}^{p}\sum_{k\neq j}{\mathbb{P}}\left(\frac{2}{n}|X_{k}^{\top}\epsilon_{j}|\geq\lambda\right)$
$\leq 2p(p-1)\exp\left(-c_{0}n\frac{\lambda^{2}}{\kappa^{4}K^{4}}\right)$
$=2\exp\left(-c_{0}n\frac{\lambda^{2}}{\kappa^{4}K^{4}}+\log p+\log(p-1)\right).$
Then, choosing $\lambda=c\sqrt{(\log p)/n}$ with $c>0$ sufficiently large ensures that the inequality $\lambda\geq(2/n)\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}$ holds for all $j$ with probability at least $1-2/p$. Hence, with this choice of $\lambda$, the inequalities (A.4) and (A.5) also hold with same probability.
Now assuming $\lambda\geq(2/n)\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}$, for the residual variances we have
$\hat{\tau}_{j}^{2}-\tau_{j}^{2}$ $=\frac{1}{n}\|X_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}-\tau_{j}^{2}=\frac{1}{n}\||X_{\setminus j}\theta_{j}+\epsilon_{j}-X_{\setminus j}\hat{\theta}_{j}\|_{2}^{2}-\tau_{j}^{2}$
$=\frac{1}{n}\|X_{\setminus j}(\hat{\theta}_{j}-\theta_{j})\|_{2}^{2}+\frac{2}{n}\langle\epsilon_{j},X_{\setminus j}(\hat{\theta}_{j}-\theta_{j})\rangle+\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}$
$\leq\frac{9}{\gamma_{\ell}}s_{j}\lambda^{2}+\frac{2}{n}\langle\epsilon_{j},X_{\setminus j}(\hat{\theta}_{j}-\theta_{j})\rangle+\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}.$
where we used (A.5). For the scalar product, Hölder’s inequality and (A.4) yield
$\frac{2}{n}|\langle\epsilon_{j},X_{\setminus j}(\hat{\theta}_{j}-\theta_{j})\rangle|\leq\frac{2}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\hat{\theta}_{j}-\theta_{j}\|_{1}\leq\frac{12}{\gamma_{\ell}}s_{j}\lambda^{2}.$
Summing up, we obtain the bound
$|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\leq\frac{21}{\gamma_{\ell}}s_{j}\lambda^{2}+\left|\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right|=c^{2}\frac{21}{\gamma_{\ell}}\frac{s_{j}\log p}{n}+\left|\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right|,$ (A.7)
provided $\lambda\geq(2/n)\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}$.
The random variable $(1/n)\|\epsilon_{j}\|_{2}^{2}$ is sub-exponential with mean $\tau_{j}^{2}$ and $\|\epsilon_{ij}\|_{\psi_{1}}\leq\kappa^{2}K^{2}$, so another application of Bernstein’s inequality yields
${\mathbb{P}}\left(\left|\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right|\geq\frac{\kappa^{2}K^{2}}{\sqrt{c_{0}/2}}\sqrt{\frac{\log p}{n}}\right)\leq 2\exp\left(-2\log p\right).$
Thus
${\mathbb{P}}\left(\left|\frac{1}{n}\|\epsilon_{j}\|_{2}^{2}-\tau_{j}^{2}\right|\geq\frac{\kappa^{2}K^{2}}{\sqrt{c_{0}/2}}\sqrt{\frac{\log p}{n}}\text{ for some }j\right)$
$\leq$ $\sum_{j=1}^{p}2\exp\left(-2\log p\right)=2p\exp(-2\log p)=2/p.$
Due to Assumption 4.1, we have $s_{j}(\log p)/n\leq M\sqrt{(\log p)/n}$, hence the inequality $|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\lesssim\sqrt{(\log p)/n}$ holds for all $j$ with probability at least $1-4/p$; this is (A.1).
Now use this bound to obtain $\hat{\tau}_{j}^{2}=\tau_{j}^{2}+O(\sqrt{(\log p)/n})$, and since $\lambda_{\mathrm{min}}(\Omega^{\star})\leq 1/\tau_{j}^{2}\leq\lambda_{\mathrm{max}}(\Omega^{\star})$ (cf. inequality (A.6)) there are $\tau_{\ell},\tau_{u}\in(0,\infty)$ such that $\tau_{\ell}\leq\hat{\tau}_{j}\leq\tau_{u}$ for all $j\in\{1,\dots,p\}$ whenever $(\log p)/n$ is small; this is (A.2).
Finally, since $\lambda=c\sqrt{(\log p)/n}\geq(2/n)\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}$ for all $j$, the regularization parameter satisfies
$\lambda\geq\max_{j}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\gamma_{u}\kappa|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\right)$
with the same probability if $c$ is sufficiently large. Likewise, because
$\sqrt{\frac{\log p}{n}}\gtrsim|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|=|\hat{\tau}_{j}+\tau_{j}||\hat{\tau}_{j}-\tau_{j}|\geq(\tau_{\ell}+\tau_{j})|\hat{\tau}_{j}-\tau_{j}|,$
it holds that
$\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|$ $=\max_{k}\left|\frac{\hat{\tau}_{j}\tau_{k}-\tau_{j}\hat{\tau}_{k}}{\hat{\tau}_{k}\tau_{k}}\right|$
$=\max_{k}\left|\frac{\left(\tau_{j}+O(\sqrt{(\log p)/n})\right)\tau_{k}-\tau_{j}\left(\tau_{k}+O(\sqrt{(\log p)/n})\right)}{\hat{\tau}_{k}\tau_{k}}\right|$
$\lesssim\sqrt{\frac{\log p}{n}}.$
Thus, if $c$ is sufficiently large then the inequality
$\lambda\geq\max_{j}\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\kappa\frac{\gamma_{u}}{\tau_{\ell}}\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|\right)$
also holds with the same probability; this is (A.3). ∎
Lemma A.2.
Under Assumptions 4.2 and 4.3, the data matrix satisfies for any vector $\nu\in{\mbox{\bf R}}^{p}$
$\lambda_{\mathrm{min}}(\Sigma)\left(1-\sqrt{\frac{p+C\log p}{n}}\right)^{2}\|\nu\|_{2}^{2}\leq\frac{1}{n}\|X\nu\|_{2}^{2}\leq\lambda_{\mathrm{max}}(\Sigma)\left(1+\sqrt{\frac{p+C\log p}{n}}\right)^{2}\|\nu\|_{2}^{2},$
with probability at least $1-2/p$, where the constant $C>0$ depends only on $\kappa$ and $K$.
Proof.
Let $\Sigma=U\Lambda U^{\top}$ be the eigenvalue decomposition of the covariance matrix, i.e., $U$ is a matrix with orthonormal columns and $\Lambda$ is a diagonal matrix. Then $X=Y\Lambda^{1/2}U^{\top}$ where $Y\in{\mbox{\bf R}}^{n\times p}$ is a matrix with i.i.d., zero mean rows that are sub-Gaussian and isotropic (have identity covariance matrix). Without loss of generality, take any unit vector $\nu\in{\mbox{\bf R}}^{p}$. Write $\tilde{\nu}=U^{\top}\nu$, which is another unit vector. We have
$\frac{1}{n}\|X\nu\|_{2}^{2}=\frac{1}{n}\|Y\Lambda^{1/2}\tilde{\nu}\|_{2}^{2}\geq\lambda_{\mathrm{min}}(\Lambda)\lambda_{\mathrm{min}}(Y^{\top}Y/n)=\lambda_{\mathrm{min}}(\Sigma)\lambda_{\mathrm{min}}(Y^{\top}Y/n).$
The upper bound
$\frac{1}{n}\|X\nu\|_{2}^{2}\leq\lambda_{\mathrm{max}}(\Sigma)\lambda_{\mathrm{max}}(Y^{\top}Y/n)$
follows similarly.
As for the sub-Gaussian norm of each row $Y_{i}$ of $Y$, we have
$\sup_{\|x\|_{2}=1}\|\langle x,X_{i}\rangle\|_{\psi_{2}}=\sup_{\|x\|_{2}=1}\|\langle x,\Sigma^{1/2}Y_{i}\rangle\|_{\psi_{2}}\geq\lambda_{\min}(\Sigma^{1/2})\sup_{\|u\|_{2}=1}\|u^{\top}Y_{i}\|_{\psi_{2}}$
and hence $\|Y_{i}\|_{\psi_{2}}\leq\sqrt{\kappa}\|X_{i}\|_{\psi_{2}}\leq\sqrt{\kappa}K$ by Assumptions 4.2 and 4.3. A straightforward application of Vershynin, (2012, Theorem 5.39) then gives the result. ∎
A.2 Proof of Theorem 4.1
Proof.
Note that under Assumption 4.1, Lemma A.2 tells us that there exist $\gamma_{\ell},\gamma_{u}\in(0,\infty)$ such that $\gamma_{\ell}\|\nu\|_{2}^{2}\leq(1/n)\|X\nu\|_{2}^{2}\leq\gamma_{u}\|\nu\|_{2}^{2}$ for all $\nu\in{\mbox{\bf R}}^{p}$ with probability $1-2/p$. Thus, by Lemma A.1, inequalities (A.1) and (A.3) hold with probability at least $1-6/p$. Throughout this proof, assume that this event occurs.
For the first part of the proof we will split the estimation error $\widehat{\Omega}-\Omega^{\star}$ as $(\widehat{\Omega}-\Omega^{\prime})+(\Omega^{\prime}-\Omega^{\star})$ and bound the the two terms separately. Here the intermediate matrix $\Omega^{\prime}$ is defined by
$\Omega_{jj}^{\prime}=1/\hat{\tau}_{j}^{2},\quad\Omega_{jk}^{\prime}=\Omega_{jk}^{\star},\quad j\neq k.$ (A.8)
For this matrix
$\lambda_{\text{min}}(\Omega^{\prime})$ $=\lambda_{\text{min}}(\Omega-\mathop{\bf diag}(\Omega-\widehat{\Omega}))$ (A.9)
$\geq\lambda_{\text{min}}(\Omega)-\max_{1\leq j\leq p}|1/\hat{\tau}_{j}^{2}-1/\tau_{j}^{2}|,$
so that by Assumption 4.3 and Lemma A.1, $\lambda_{\text{min}}(\Omega^{\prime})\geq 0$ whenever $(\log p)/n$ is sufficiently small. This means that $\Omega^{\prime}$ is feasible in the joint partial regression problem (3.1).
To alleviate the notation, let $\omega^{\star}_{j}=\Omega_{\setminus j,j}^{\star}$. Define the shifted variables $\nu_{j}=\Omega_{\setminus j,j}-\omega^{\star}_{j}$, and define $G$ as the objective function of (3.1) in these variables, i.e.,
$G(\nu_{1},\dots,\nu_{p})=\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}(\omega^{\star}_{j}+\nu_{j})\|_{2}^{2}+\lambda\hat{\tau}^{2}_{j}\|\omega^{\star}_{j}+\nu_{j}\|_{1}\right).$
Now recall that $X_{j}=-\tau_{j}^{2}X_{\setminus j}\omega^{\star}_{j}+\epsilon_{j}$. Hence we can write
$G(\nu_{1},\dots,\nu_{p})$ $=\sum_{j=1}^{p}\left(\frac{1}{2n}\|\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}\nu_{j}\|_{2}^{2}+\lambda\hat{\tau}_{j}^{2}\|\omega^{\star}_{j}+\nu_{j}\|_{1}\right)$ (A.10)
$=\sum_{j=1}^{p}\bigg{(}\frac{1}{2n}\|\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j}\|_{2}^{2}+\frac{\hat{\tau}_{j}^{2}}{n}\langle\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j},X_{\setminus j}\hat{\nu}_{j}\rangle$
$\qquad+\frac{1}{2n}\|\hat{\tau}_{j}^{2}X_{\setminus j}\nu_{j}\|_{2}^{2}+\lambda\hat{\tau}_{j}^{2}\|\omega^{\star}_{j}+\nu_{j}\|_{1}\bigg{)}$
In particular,
$G(0,\dots,0)=\sum_{j=1}^{p}\left(\frac{1}{2n}\|\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j}\|_{2}^{2}+\lambda\hat{\tau}_{j}^{2}\|\omega^{\star}_{j}\|_{1}\right).$ (A.11)
Put $\hat{\nu}_{j}=\widehat{\Omega}_{\setminus j,j}-\omega^{\star}_{j}$. Then by construction we have the inequality
$G(\hat{\nu}_{1},\dots,\hat{\nu}_{p})\leq G(0,\dots,0).$ (A.12)
Insert (A.10) with $\nu_{j}=\hat{\nu}_{j}$ and (A.11) into (A.12), rearrange and use $(1/n)\|X\nu\|_{2}^{2}\geq\gamma_{\ell}\|\nu\|_{2}^{2}$ to obtain
$\sum_{j=1}^{p}$ $\frac{1}{2}\gamma_{\ell}\hat{\tau}_{j}^{4}\|\hat{\nu}_{j}\|_{2}^{2}$ (A.13)
$\leq\sum_{j=1}^{p}\left(\lambda\hat{\tau}_{j}^{2}(\|\omega^{\star}_{j}\|_{1}-\|\omega^{\star}_{j}+\hat{\nu}_{j}\|_{1})-\frac{\hat{\tau}_{j}^{2}}{n}\langle\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})X_{\setminus j}\omega^{\star}_{j},X_{\setminus j}\hat{\nu}_{j}\rangle\right).$
Starting with the scalar product term on the right-hand side of this inequality, Hölder’s inequality, $\lambda_{\text{max}}((1/n)X^{\top}X)\leq\gamma_{u}$, $\lambda_{\text{min}}(\Sigma)\geq 1/\kappa$ and finally (A.3) yield
$\bigg{|}\frac{1}{n}\langle\epsilon_{j}+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})$ $X_{\setminus j}\omega^{\star}_{j},X_{\setminus j}\hat{\nu}_{j}\rangle\bigg{|}=\bigg{|}\frac{1}{n}\langle\epsilon_{j},X_{\setminus j}\hat{\nu}_{j}\rangle+(\hat{\tau}_{j}^{2}-\tau_{j}^{2})\frac{1}{n}\langle X_{\setminus j}\omega^{\star}_{j},X_{\setminus j}\hat{\nu}_{j}\rangle\bigg{|}$ (A.14)
$\leq\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\hat{\nu}_{j}\|_{1}+|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\frac{1}{n}\|X_{\setminus j}^{\top}X_{\setminus j}\|_{2}\|\omega^{\star}_{j}\|_{2}\|\hat{\nu}_{j}\|_{2}$
$\leq\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\hat{\nu}_{j}\|_{1}+\gamma_{u}\kappa|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\|\hat{\nu}_{j}\|_{1}$
$=\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\gamma_{u}\kappa|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\right)\|\hat{\nu}_{j}\|_{1}$
$\leq\lambda\|\hat{\nu}_{j}\|_{1}.$
Now consider the norm terms in (A.13). Let $S_{j}$ be the indices of the true non-zero elements of $\omega^{\star}_{j}$, and $S_{j}^{c}$ be the indices of the zero elements. Then, using the reverse triangle ineqality we have
$\|\omega^{\star}_{j}+\hat{\nu}_{j}\|_{1}=\|(\omega^{\star}_{j}+\hat{\nu}_{j})_{S_{j}}\|_{1}+\|(\hat{\nu}_{j})_{S_{j}^{c}}\|_{1}\geq\|\omega^{\star}_{j}\|_{1}-\|(\hat{\nu}_{j})_{S_{j}}\|_{1}+\|(\hat{\nu}_{j})_{S_{j}^{c}}\|_{1},$
and hence
$\|\omega^{\star}_{j}\|_{1}-\|\omega^{\star}_{j}+\hat{\nu}_{j}\|_{1}\leq\|(\hat{\nu}_{j})_{S_{j}}\|_{1}-\|(\hat{\nu}_{j})_{S_{j}^{c}}\|_{1}.$ (A.15)
Inequalities (A.14) and (A.15) together with (A.13) show that
$\sum_{j=1}^{p}\frac{1}{2}\gamma_{\ell}\hat{\tau}_{j}^{4}\|\hat{\nu}_{j}\|_{2}^{2}\leq\sum_{j=1}^{p}\lambda\hat{\tau}_{j}^{2}(\|(\hat{\nu}_{j})_{S_{j}}\|_{1}-\|(\hat{\nu}_{j})_{S_{j}^{c}}\|_{1}+\|\hat{\nu}_{j}\|_{1})=\sum_{j=1}^{p}2\lambda\hat{\tau}_{j}^{2}\|(\hat{\nu}_{j})_{S_{j}}\|_{1}.$
Interpreting the sums over columns as matrix norms we have
$\frac{1}{2}\gamma_{\ell}\min_{j}\hat{\tau}_{j}^{4}\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F}^{2}\leq 2\max_{j}\hat{\tau}_{j}^{2}\lambda\|(\widehat{\Omega}-\Omega^{\prime})_{S}\|_{1}\leq 2\max_{j}\hat{\tau}_{j}^{2}\lambda\sqrt{s}\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F},$
where $S$ are the indices of the true non-zero entries, and thus
$\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F}\leq\frac{4}{\gamma_{\ell}}\frac{\max_{j}\hat{\tau}^{2}_{j}}{\min_{j}\hat{\tau}^{4}_{j}}\sqrt{s}\lambda.$ (A.16)
By (A.1), $\hat{\tau}_{j}^{2}=\tau_{j}^{2}+O(\sqrt{(\log p)/n})$, so since $\lambda_{\mathrm{min}}(\Omega^{\star})\leq 1/\tau_{j}^{2}\leq\lambda_{\mathrm{max}}(\Omega^{\star})$ there exist $\tau_{\ell},\tau_{u}\in(0,\infty)$ such that $\tau_{\ell}\leq\hat{\tau}_{j}\leq\tau_{u}$ for all $j\in\{1,\dots,p\}$ whenever $(\log p)/n$ is small. Hence $\max_{j}\hat{\tau}^{2}_{j}/\min_{j}\hat{\tau}^{4}_{j}$ is bounded, so (A.16) implies that $\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F}\lesssim\sqrt{s(\log p)/n}$.
It now remains to show that $\Omega^{\prime}-\Omega^{\star}$ is small. Recall (Equation (A.8)) that these matrices agree on the off-diagonal. On the diagonal we have
$|\Omega^{\prime}_{jj}-\Omega^{\star}_{jj}|=|1/\hat{\tau}_{j}^{2}-1/\tau_{j}^{2}|\leq\frac{1}{\hat{\tau}_{j}^{2}\tau_{j}^{2}}|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\leq\frac{\lambda_{\mathrm{max}}(\Omega^{\star})}{\tau_{\ell}^{2}}|\hat{\tau}_{j}^{2}-\tau_{j}^{2}|\lesssim\sqrt{\frac{\log p}{n}}.$
Hence
$\|\widehat{\Omega}-\Omega^{\star}\|_{\rm F}^{2}=\|\widehat{\Omega}-\Omega^{\prime}\|_{\rm F}^{2}+\sum_{j=1}^{p}(\Omega^{\prime}_{jj}-\Omega^{\star}_{jj})^{2}\lesssim\frac{(s+p)\log p}{n},$
which is the stated result for the precision matrix.
We now turn to the partial correlation matrix. Similarly to as above, write $q_{j}^{\star}=Q_{\setminus j,j}^{\star}$ and define the variables $\nu_{j}=Q_{\setminus j,j}-q_{j}^{\star}$ as well as the function
$H(\nu_{1},\dots,\nu_{p})=\sum_{j=1}^{p}\left(\frac{1}{2n}\|X_{j}-X_{\setminus j}\widehat{R}_{j}(q^{\star}_{j}+\nu_{j})\|_{2}^{2}+\lambda\|\widehat{R}_{j}(q^{\star}_{j}+\nu_{j})\|_{1}\right),$
with $\widehat{R}_{j}=\mathop{\bf diag}(\hat{\tau}_{j}/\hat{\tau}_{1},\dots,\hat{\tau}_{j}/\hat{\tau}_{j-1},\hat{\tau}_{j}/\hat{\tau}_{j+1},\dots,\hat{\tau}_{j}/\hat{\tau}_{p})$.
Proceeding like previously, we note that $X_{j}=\tau_{j}^{2}X_{\setminus j}Q_{\setminus j,j}^{\star}+\epsilon_{j}=X_{\setminus j}R_{j}q_{j}^{\ast}+\epsilon_{j}$ with $R_{j}=\mathop{\bf diag}(\tau_{j}/\tau_{1},\dots,\tau_{j}/\tau_{j-1},\tau_{j}/\tau_{j+1},\dots,\tau_{j}/\tau_{p})$, and then
$H(\nu_{1},\dots,\nu_{p})$ $=\sum_{j=1}^{p}\left(\frac{1}{2n}\|\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j}-X_{\setminus j}\widehat{R}_{j}\nu_{j}\|_{2}^{2}+\lambda\|\widehat{R}_{j}(q^{\star}_{j}+\nu_{j})\|_{1}\right)$
$=\sum_{j=1}^{p}\bigg{(}\frac{1}{2n}\|\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j}\|_{2}^{2}-\frac{1}{n}\langle\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j},X_{\setminus j}\widehat{R}_{j}\nu_{j}\rangle$
$\qquad+\frac{1}{2n}\|X_{\setminus j}\widehat{R}_{j}\nu_{j}\|_{2}^{2}+\lambda\|\widehat{R}_{j}(q^{\star}_{j}+\nu_{j})\|_{1}\bigg{)}.$
In particular,
$H(0,\dots,0)=\sum_{j=1}^{p}\left(\frac{1}{2n}\|\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j}\|_{2}^{2}+\lambda\|\widehat{R}_{j}q^{\star}_{j}\|_{1}\right).$
The true partial correlation matrix $Q^{\star}$ is always feasible in the joint partial regression problem, so for $\tilde{\nu}_{j}=\widehat{Q}_{\setminus j,j}-q_{j}^{\star}$ we have by definition $H(\tilde{\nu}_{1},\dots,\tilde{\nu}_{p})\leq H(0,\dots,0)$. Inserting the above expressions for $H(\nu_{1},\dots,\nu_{p})$ with $\nu_{j}=\tilde{\nu}_{j}$ and $H(0,\dots,0)$ respectively into this inequality and rearranging terms, we arrive at
$\sum_{j=1}^{p}\frac{\gamma_{\ell}}{2}$ $\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{2}^{2}$
$\leq\sum_{j=1}^{p}\Big{(}\lambda(\|\widehat{R}_{j}q^{\star}_{j}\|_{1}-\|\widehat{R}_{j}(q^{\star}_{j}+\tilde{\nu}_{j})\|_{1})+\frac{1}{n}\langle\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j},X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\rangle\Big{)}.$
Once again using Hölder’s inequality we have
$\bigg{|}\frac{1}{n}\langle\epsilon_{j}-X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j},$ $X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\rangle\bigg{|}=\left|\frac{1}{n}\langle\epsilon_{j},X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\rangle-\frac{1}{n}\langle X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j},X_{\setminus j}\widehat{R}_{j}\tilde{\nu}_{j}\rangle\right|$
$\leq\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{1}+\frac{1}{n}\|X_{\setminus j}^{\top}X_{\setminus j}(\widehat{R}_{j}-R_{j})q^{\star}_{j}\|_{2}\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{2}$
$\leq\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{1}+\gamma_{u}\lambda_{\mathrm{max}}(\widehat{R}_{j}-R_{j})\|q^{\star}_{j}\|_{2}\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{1}$
$\leq\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{1}+\kappa\frac{\gamma_{u}}{\tau_{\ell}}\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{1}$
$=\left(\frac{1}{n}\|X_{\setminus j}^{\top}\epsilon_{j}\|_{\infty}+\kappa\frac{\gamma_{u}}{\tau_{\ell}}\max_{k}\left|\frac{\hat{\tau}_{j}}{\hat{\tau}_{k}}-\frac{\tau_{j}}{\tau_{k}}\right|\right)\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{1}.$
The bracketed expression on the right-hand side above is bounded by (A.3), so that
$\sum_{j=1}^{p}\frac{\gamma_{\ell}}{2}\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{2}^{2}\leq\lambda\sum_{j=1}^{p}(\|\widehat{R}_{j}q^{\star}_{j}\|_{1}-\|\widehat{R}_{j}q^{\star}_{j}+\widehat{R}_{j}\tilde{\nu}_{j}\|_{1}+\|\widehat{R}_{j}\tilde{\nu}_{j}\|_{1})$
Recall that $\widehat{R}_{j}$ is diagonal, bound the left-hand side above from below and adapt the inequality (A.15) for the right-hand side to conclude that
$\sum_{j=1}^{p}\frac{\gamma_{\ell}}{2}\left(\frac{\tau_{\ell}}{\tau_{u}}\right)^{2}\|\tilde{\nu}_{j}\|_{2}^{2}\leq 2\lambda\sum_{j=1}^{p}\|(\widehat{R}_{j}\tilde{\nu}_{j})_{S_{j}}\|_{1}\leq 2\lambda\frac{\tau_{u}}{\tau_{\ell}}\sum_{j=1}^{p}\|(\tilde{\nu}_{j})_{S_{j}}\|_{1}.$
Rearranging terms and using $\sum_{j}\|\tilde{\nu}_{j}\|_{2}^{2}=\|\widehat{Q}-Q^{\star}\|_{\rm F}^{2}$ and $\sum_{j}\|\tilde{\nu}_{j}\|_{1}=\|\widehat{Q}-Q^{\star}\|_{1}$, we obtain
$\|\widehat{Q}-Q^{\star}\|_{\rm F}^{2}\leq\frac{4}{\gamma_{\ell}}\left(\frac{\tau_{u}}{\tau_{\ell}}\right)^{3}\lambda\|(\widehat{Q}-Q^{\star})_{S}\|_{1}\leq\frac{4}{\gamma_{\ell}}\left(\frac{\tau_{u}}{\tau_{\ell}}\right)^{3}\lambda\sqrt{s}\|\widehat{Q}-Q^{\star}\|_{\rm F}.$
Inserting $\lambda=c\sqrt{(\log p)/n}$ and dividing by the Frobenius error, we arrive at the bound
$\|\widehat{Q}-Q^{\star}\|_{\rm F}\leq\frac{4c}{\gamma_{\ell}}\left(\frac{\tau_{u}}{\tau_{\ell}}\right)^{3}\sqrt{\frac{s\log p}{n}}.$
This completes the proof. ∎
Appendix B Computational details
Primal update.
The primal iterate update is a projection onto the set $\mathcal{S}$,
$\Omega^{(k+1)}=\Pi_{\mathcal{S}}\left(\Omega^{(k)}-\gamma U^{(k)}-\gamma\nabla f(\Omega^{(k)})\right).$
Importantly, note that the matrix which is being projected is not in general symmetric. The projection of a matrix $A\in{\mbox{\bf R}}^{n\times n}$ onto $\mathcal{S}$ is by definition given by
$\Pi_{\mathcal{S}}(A)=\operatornamewithlimits{argmin}_{\Omega\in\mathcal{S}}\|\Omega-A\|_{\rm F}^{2}.$
Due to symmetry, for any $\Omega\in\mathcal{S}$ the squared distance $\|\Omega-A\|_{\rm F}^{2}$ can be written as
$\sum_{j=1}^{p}(\Omega_{jj}-A_{jj})^{2}+2\sum_{j=1}^{p}\sum_{k<j}^{p}(\Omega_{jk}-(A_{jk}+A_{kj})/2)^{2}=\|\Omega-(A+A^{\top})/2\|_{\rm F}^{2}$ (B.1)
up to a term constant in $\Omega$. Let $Q\Lambda Q^{\top}=(A+A^{\top})/2$ be the eigenvalue decomposition of the symmetric part of $A$. The Frobenius norm is invariant under orthogonal transformations, so (B.1) can equivalently be written as
$\|Q^{\top}(\Omega-(A+A^{\top})/2)Q\|_{\rm F}^{2}=\|Q^{\top}\Omega Q-\Lambda\|_{\rm F}^{2}.$
Clearly, this expression is minimized by choosing $\Omega$ such that $Q^{\top}\Omega Q$ is diagonal, with the diagonal elements being the projection of the eigenvalues of ($A+A^{\top})/2$ onto the interval $[\alpha,\beta]$. Thus the projection $\Pi_{\mathcal{S}}(A)$ is given by projecting the symmetric part of $A$,
$\Pi_{\mathcal{S}}(A)$ $=Q\Lambda_{[\alpha,\beta]}Q^{\top}$
where $\Lambda_{[\alpha,\beta]}$ is a diagonal matrix with entries $(\Lambda_{[\alpha,\beta]})_{jj}=\min(\max(\Lambda_{jj},\alpha),\beta)$.
The gradient that appears in the primal step is given by
$\nabla f(\Omega)_{jj}$ $=0$
$\nabla f(\Omega)_{\setminus j,j}$ $=\hat{\tau}_{j}^{2}X_{\setminus j}^{\top}\nabla\ell(X_{j}+\hat{\tau}_{j}^{2}X_{\setminus j}\Omega_{\setminus j,j}),$
and depends on the loss function we choose. If we take the quadratic loss as in (3.1), then for $i=1,\dots,n$,
$\nabla\ell(z)_{i}=\frac{z_{i}}{n}.$
For the Huber loss
$\ell(z)=\frac{1}{n}\sum_{i=1}^{n}\phi_{\rho}(z_{i}),\quad\phi_{\rho}(r)=\begin{cases}r^{2}/2,&\text{if }|r|\leq\rho,\\ \rho(|r|-\rho/2),&\text{if }|r|>\rho,\end{cases}$
we have instead
$\nabla\ell(z)_{i}=\frac{1}{n}\begin{cases}z_{i},&\text{if }|z_{i}|\leq\rho,\\ \rho\mathop{\bf sign}(z_{i}),&\text{otherwise}.\end{cases}$
For two points $\Omega$ and $\widetilde{\Omega}$ in ${\mbox{\bf R}}^{n\times n}$, with either the quadratic or Huber loss, $f$ satisfies
$\|\nabla f(\Omega)-\nabla f(\Omega^{\prime})\|_{\rm F}$ $\leq\frac{1}{n}\left(\sum_{j=1}^{p}\|\hat{\tau}_{j}^{4}X_{\setminus j}^{\top}X_{\setminus j}(\Omega_{\setminus j,j}-\Omega^{\prime}_{\setminus j,j})\|_{2}^{2}\right)^{1/2}$
$\leq\frac{1}{n}\max_{1\leq j\leq p}\hat{\tau}_{j}^{4}\lambda_{\text{max}}(X_{\setminus j}^{\top}X_{\setminus j})\|\Omega-\Omega^{\prime}\|_{\rm F}.$
So $f$ is $L$-smooth with $L=\max_{1\leq j\leq p}\hat{\tau}_{j}^{4}\lambda_{\text{max}}(X_{\setminus j}^{\top}X_{\setminus j})/n$.
Dual update.
By the Moreau identity, we can write the dual iterate update with the proximal operator of $g/\eta$ instead of the Fenchel conjugate $\eta g^{*}$,
$V^{(k+1)}$ $=U^{(k)}+\eta(2\Omega^{(k+1)}-\Omega^{(k)})+\gamma\eta(\nabla f(\Omega^{(k)})-\nabla f(\Omega^{(k+1)})),$
$U^{(k+1)}$ $=V^{(k+1)}-\eta\mathbf{Prox}_{g/\eta}(V^{(k+1)}/\eta).$
The proximal operator of $g/\eta$ is given by
$\mathbf{Prox}_{g/\eta}(V^{(k+1)}/\eta)_{jj}$ $=1/\hat{\tau}_{j}^{2},$
$\mathbf{Prox}_{g/\eta}(V^{(k+1)}/\eta)_{kj}$ $=\mathop{\bf sign}(V_{kj}^{(k+1)})(V_{kj}^{(k+1)}-\hat{\tau}_{j}^{2}\lambda_{j})_{+}/\eta,$
for $j=1,\dots,p$ and $k\neq j$.